### v2

# DELIFT: Data Efficient Language model Instruction Fine-Tuning

Abstract
Fine-tuning large language models (LLMs) is essential for enhancing their performance on specific tasks but is often resource-intensive due to redundant or uninformative data. To address this inefficiency, we introduce DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel algorithm that systematically optimizes data selection across the three key stages of fine-tuning: (1) instruction tuning, (2) task-specific fine-tuning (e.g., reasoning, question-answering), and (3) continual fine-tuning (e.g., incorporating new data versions). Unlike existing methods that focus on single-stage optimization or rely on computationally intensive gradient calculations, DELIFT operates efficiently across all stages. Central to our approach is a pairwise utility metric that quantifies how beneficial a data sample is for improving the model’s responses to other samples, effectively measuring the informational value relative to the model’s current capabilities. By leveraging different submodular functions applied to this metric, DELIFT selects diverse and optimal subsets that are useful across all stages of fine-tuning. Experiments across various tasks and model scales demonstrate that DELIFT can reduce the fine-tuning data size by up to 70% without compromising performance, offering significant computational savings and outperforming existing methods in both efficiency and efficacy.
1 Introduction
Fine-tuning large language models (LLMs) is pivotal for adapting these powerful architectures (Devlin et al., 2019; Brown et al., 2020a; Touvron et al., 2023) to specialized tasks such as intricate reasoning, precise question-answering, and the seamless integration of new information (Ouyang et al., 2022). This transformation—from a general-purpose model to a task-specific agent—heavily relies on the quality and nature of the data employed during fine-tuning, which critically determines the model’s subsequent performance (Wei et al., 2022; Zhou et al., 2023; Hoffmann et al., 2024).
The effectiveness of fine-tuning hinges on the quality, diversity, and relevance of the selected data (Gururangan et al., 2020; Wei et al., 2022; Zhou et al., 2023). High-quality data ensures accurate learning, diverse data enhances generalization, and relevant data aligns the model’s capabilities with specific application needs. However, optimizing data selection across different fine-tuning phases remains a significant challenge, leading to our central research question:
How can we create a unified framework for efficient data selection across all fine-tuning stages of LLMs, while optimizing performance and maximizing data efficiency?
To address this challenge, we present DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel, unified, and computationally efficient algorithm engineered to optimize data selection across all stages of the fine-tuning process. The key innovation of DELIFT lies in its pairwise utility metric, which assesses the informational value of data samples relative to both the model’s current capabilities and other samples within the dataset. This metric, combined with submodular optimization techniques, allows DELIFT to efficiently select optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs.
The typical fine-tuning process comprises three key stages: 1. Instruction Tuning: Enhances the model’s ability to follow general instructions (Mishra et al., 2022; Wei et al., 2022; Longpre et al., 2023); 2. Task-Specific Fine-Tuning: Refines the model’s expertise in specific domains (Gururangan et al., 2020; Cobbe et al., 2021); 3. Continual Fine-tuning: Enables the model to integrate new information while mitigating catastrophic forgetting (Madotto et al., 2021; Wu et al., 2024). DELIFT is able to optimize data selection processes across all three stages. Additionally, DELIFT offers significant benefits for In-Context Learning (ICL) (Brown et al., 2020b; Xue et al., 2024). By utilizing the selected subsets as the ICL example pool, DELIFT achieves similar or better performance compared to using the entire dataset, thereby enhancing data efficiency in ICL scenarios. This dual functionality is empirically validated in our experimental results.
Existing data selection methodologies often fail to address the nuanced requirements of the aforementioned distinct fine-tuning stages. Many approaches are tailored to a single stage, lacking the adaptability needed for comprehensive fine-tuning (Xia et al., 2024; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024). Others depend on computationally intensive procedures, such as exhaustive gradient computations, rendering them impractical for large-scale models and datasets (Killamsetty et al., 2021b; a; Xia et al., 2024; Zhang et al., 2024). Additionally, some methods utilize features obtained from an independent model that are not specifically aligned with the model undergoing fine-tuning, reducing their effectiveness (Killamsetty et al., 2023; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024; Du et al., 2023).
DELIFT addresses these limitations by adapting to the unique requirements of each fine-tuning stage. 1. Instruction Tuning: Selects diverse data to enhance general instruction-following capabilities; 2. Task-Specific Fine-Tuning: Prioritizes data that is aligned with the target task, to refine specialized expertise; 3. Continual Fine-tuning: Identifies novel, complementary information to expand the model’s knowledge base while safeguarding against catastrophic forgetting.
Figure 1 illustrates how DELIFT optimizes data selection across these stages, demonstrating the selection and pruning processes in each fine-tuning phase. By leveraging submodular optimization techniques (Fujishige, 2005; Bilmes, 2022) and submodular information measures (Iyer et al., 2021), DELIFT efficiently selects optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs. This approach effectively balances data utility and computational efficiency.
Figure 1: DELIFT data selection across fine-tuning stages. (a) Instruction Tuning: Diverse instructions selected; redundant samples pruned. (b) Task-Specific Fine-Tuning: Mutually informative (with benchmark data) and diverse samples are prioritized for selection. (c) Continual Fine-tuning: New samples that are novel are integrated; new samples with overlapping information are pruned.
a)Refer to caption b)Refer to caption c)Refer to caption
Our key contributions are as follows:
1) Versatile Pairwise Utility Metric: A novel, easy-to-compute metric for assessing data informativeness, incorporating model feedback applicable across all fine-tuning stages.
2) Unified Data Selection Algorithm: DELIFT systematically optimizes data selection for instruction tuning, task-specific fine-tuning, and continual fine-tuning within a single framework.
3) Computational Efficiency: Circumvents resource-intensive operations, ensuring scalability to large datasets and models. DELIFT achieves at least 70% reduction in computational time compared to gradient-based methods on benchmark tasks.
4) Enhanced Performance with Reduced Data: Demonstrates the ability to reduce fine-tuning data size by up to 70% without compromising performance, and achieves comparable efficacy as to utilizing the full dataset.
5) Improvement over Existing Methods: Outperforms current data selection techniques by up to 26% in effectiveness across diverse tasks and model scales (see Section  4).
The remainder of this paper is organized as follows: Section 2 provides background on fine-tuning LLMs and reviews related work. Section 3 details the methodology behind DELIFT, including the development of our pairwise utility metric and the submodular optimization process. Section 4 presents experimental results that showcase the effectiveness and efficiency of our method. Section 5 discusses the implications of our findings and potential future directions. Finally, we release our code base for further research.
2 Related Work
Efficient data subset selection is vital for enhancing training efficiency in deep neural networks while maintaining or improving model performance. This section categorizes existing subset selection methods into model-independent and model-dependent approaches and identifies the gaps our work addresses. Model-independent subset selection methods focus on selecting representative subsets without model-specific feedback. Common approaches include using pre-trained sentence embeddings with distance or clustering metrics (Bukharin & Zhao, 2024; Sorscher et al., 2023; Killamsetty et al., 2023; Du et al., 2023; Bhatt et al., 2024), as well as employing large models like GPT-4 or pre-trained reward models for high-quality data filtering (Du et al., 2023; Chen et al., 2024). However, these methods often struggle to translate the assessed diversity or quality into downstream utility. Model-dependent subset selection aims to identify data samples beneficial to the downstream model by analyzing features like per-sample gradients or loss values. Methods such as GradMatch (Killamsetty et al., 2021a), CRAIG (Mirzasoleiman et al., 2020), and TAGCOS (Zhang et al., 2024) focus on selecting samples that approximate the gradient updates over the full dataset. GLISTER (Killamsetty et al., 2021b) employs bilevel optimization to align gradients from selected subsets with those of a validation set. LESS (Xia et al., 2024) proposes computing gradients through LoRA fine-tuning to reduce the computational cost of gradient computation and utilizes random projection to address gradient dimensionality issues. Li et al. (2024) proposed the IFD score, a computationally efficient model-dependent metric that assesses instruction difficulty to filter challenging samples, though it does not guarantee data diversity. While effective in capturing useful samples, these methods often face computational challenges, especially with LLMs. Persistent limitations across these methods include: (i) Limited Adaptability across different fine-tuning stages, (ii) Computational Intensity due to model feedback reliance, (iii) Lack of Unified Solutions applicable across all fine-tuning phases, and (iv) ineffective Redundancy Handling. DELIFT addresses these limitations through a novel pairwise utility metric, which effectively aligns with the model’s evolving capabilities throughout fine-tuning. By integrating submodular optimization with pairwise model-dependent metrics that evaluate relative sample utility, DELIFT minimizes redundancy while maximizing adaptability and computational efficiency. This approach proves effective across diverse use cases including instruction tuning, task-specific fine-tuning, continual fine-tuning, and In-Context Learning (ICL), offering a versatile and scalable solution for data subset selection.
3 Methodology
This section presents foundational concepts and the specific approach of DELIFT, focusing on data subset selection through a utility-based kernel integrated with submodular optimization techniques.
3.1 Notation
Let $\mathcal{D}$ denote the fine-tuning dataset, comprising elements $d_{i}=(x_{i},y_{i})$, where $x_{i}$ is the input sequence and $y_{i}$ is the corresponding output sequence. Our objective is to select a subset $\mathcal{A}\subseteq\mathcal{D}$ that maximizes the model’s performance while minimizing computational costs. The selection strategy adapts based on the fine-tuning objective, which may include instruction tuning, task-specific adaptation, or continual learning.
3.2 Utility-Based Kernel
At the core of DELIFT lies the utility-based kernel, a mechanism designed to quantify the informativeness of one data point when used as an in-context example for another. Consider two data points, $(x_{i},y_{i})$ and $(x_{j},y_{j})$. The utility of data point $j$ relative to data point $i$, denoted as $UF_{ij}$, is defined as:
$UF_{ij}=d(GT_{i},p(y_{i}\mid x_{i}))-d(GT_{i},p(y_{i}\mid x_{j},y_{j},x_{i})),$ (1)
where: $d(\cdot,\cdot)$ is a length-normalized distance metric between two probability distributions, $GT_{i}$ is the ground truth distribution for the sequence $y_{i}$, modeled as a vector of ones for each token to signify perfect prediction, $p(y_{i}\mid x_{i})$ is the model’s predicted probability distribution for $y_{i}$ given only the input $x_{i}$, $p(y_{i}\mid x_{j},y_{j},x_{i})$ is the predicted distribution for $y_{i}$ when the model is provided with $(x_{j},y_{j})$ as an in-context example followed by $x_{i}$.
The distance metric $d(p,q)$ is calculated using the length normalized L2 norm and is defined as:
$d(p,q)=\sqrt{\frac{\sum_{k=1}^{N}(p_{k}-q_{k})^{2}}{N}},$ (2)
where $p_{k}$ and $q_{k}$ are the $k$-th elements of the flattened probability distributions $p$ and $q$, respectively. Importantly, $N$ varies with each data sample and corresponds to the number of tokens in the ground truth response $y_{i}$. This normalization ensures that the distance measure remains scale-invariant across different sequence lengths and vocabulary sizes.
To compute the probability distributions accurately, we employ the teacher forcing technique (Williams & Zipser, 1989). This method ensures that the model uses the ground truth of previous tokens when predicting each subsequent token in the sequence, enabling reliable measurement of prediction accuracy.
The utility value $UF_{ij}$ measures the improvement in prediction accuracy for the entire sequence $(x_{i},y_{i})$ when utilizing $(x_{j},y_{j})$ as an in-context example. A positive $UF_{ij}$ indicates that including data point $j$ enhances the model’s prediction accuracy for $i$, whereas a negative value suggests an adverse effect.
3.3 Submodular Functions for Dataset Selection
To optimize the selection of informative data subsets, DELIFT leverages submodular functions (Fujishige, 2005). Submodular functions are characterized by the property of diminishing marginal returns, making them ideal for selecting diverse, informative, and non-redundant subsets. Submodular function maximization can be efficiently approximated using a greedy algorithm, with a provable approximation guarantee of $1-\frac{1}{e}$ of the optimal solution (Nemhauser et al., 1978).
We employ three tailored submodular functions (Iyer et al., 2021), each suited to a specific fine-tuning stage:
3.3.1 Facility Location (FL)
From an information perspective, the Facility Location function maximizes the coverage of the information space. It ensures that the selected subset $\mathcal{A}$ contains examples that are collectively representative of the entire dataset’s $\mathcal{D}$ information content. This is particularly useful in instruction tuning, where we aim to capture a diverse range of instruction types and their informational content. It is defined as the following where $s_{ij}$ is the similarity measure between data points $i$ and $j$:
$f_{FL}(\mathcal{A})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij},$ (3)
3.3.2 Facility Location Mutual Information (FLMI)
The FLMI function is designed to maximize the mutual information between the selected subset $\mathcal{A}$ and the target domain dataset $\mathcal{D_{T}}$. In our context, it ensures that the selected data points are not only informative in general but also particularly relevant to the specific task at hand. This makes it ideal for task-specific fine-tuning, where we want to bridge the gap between general knowledge and task-specific information. It is defined below where $\eta$ is a scaling factor (set to $1$ in our experiments):
$f_{FLMI}(\mathcal{A};\mathcal{D}_{T})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}+\eta\sum_{j\in\mathcal{A}}\max_{i\in\mathcal{D}_{T}}s_{ij},$ (4)
3.3.3 Facility Location Conditional Gain (FLCG)
From an information-theoretic standpoint, the FLCG function aims to maximize the conditional information gain of the selected subset $\mathcal{A}$ given the existing dataset $\mathcal{D_{E}}$. It quantifies how much new information each data point brings, conditional on what the model already knows. This is crucial for continual fine-tuning, where we want to avoid redundancy and focus on novel, complementary information that expands the model’s knowledge base without unnecessary repetition. It is defined as the following where $\nu$ is a scaling factor (set to $1$ in our experiments).
$f_{FLCG}(\mathcal{A}\mid\mathcal{D}_{E})=\sum_{i\in\mathcal{D}}\max\left(\max_{j\in\mathcal{A}}s_{ij}-\nu\max_{k\in\mathcal{D}_{E}}s_{ik},0\right),$ (5)
Each submodular function, when combined with our utility-based kernel, guides the selection of data subsets tailored to the specific fine-tuning stage. This ensures that DELIFT selects the most informative and diverse examples, maximizing the efficiency and effectiveness of fine-tuning.
3.4 Utility Kernel as Feature Space
Our approach utilizes the utility-based kernel as a feature space for data selection, representing a significant departure from traditional semantic similarity-based methods. Traditional methods often rely on sentence embeddings (SE) to capture static semantic similarities between data points. In contrast, our utility-based kernel measures the actual impact of examples on model performance, providing a dynamic and task-specific assessment.
This distinction is crucial for two main reasons: 1. Semantic Diversity vs. Performance Enhancement: While SE-based methods select diverse examples solely based on semantic content, our utility-based approach selects examples that demonstrably improve model performance across various inputs; 2. Model-Aware Selection: The utility-based kernel is attuned to the model’s current capabilities and weaknesses, enabling the selection of data points that are most beneficial for enhancing performance on the target task.
By integrating the utility-based kernel with the aforementioned submodular functions DELIFT tailors the data selection process to each fine-tuning stage: instruction tuning, task-specific fine-tuning, and continual learning.
3.5 Data Subset Selection Algorithm
To operationalize our data selection approach, we employ a **greedy algorithm** that iteratively builds the subset $\mathcal{A}$ by selecting the data point that offers the maximum marginal gain in the chosen submodular function.
Algorithm 1 Greedy Maximization for Submodular Function
1:Dataset $\mathcal{D}$, submodular function $f$, budget $k$
2:Initialize subset $\mathcal{A}\leftarrow\emptyset$
3:for $t=1$ to $k$ do
4:     Select $d^{*}=\arg\max_{d\in\mathcal{D}\setminus\mathcal{A}}\left(f(\mathcal{A}\cup\{d\})-f(\mathcal{A})\right)$
5:     Update $\mathcal{A}\leftarrow\mathcal{A}\cup\{d^{*}\}$
6:end for
7:return $\mathcal{A}$
This greedy algorithm ensures that each addition to the subset $\mathcal{A}$ maximizes the marginal gain in the submodular function $f$. By iteratively selecting the most beneficial data points according to the utility-based kernel and the specific submodular function tailored to the fine-tuning stage, DELIFT efficiently utilizes the data budget to select the most informative examples.
The complete subset selection process involves the following steps: 1. Compute the Utility-Based Kernel: Calculate $UF_{ij}$ for all relevant pairs of data points in the dataset to assess their informativeness; 2. Select the Appropriate Submodular Function: Depending on the fine-tuning stage (instruction tuning, task-specific fine-tuning, or continual fine-tuning), choose the corresponding submodular function (FL, FLMI, or FLCG); 3. Apply the Greedy Maximization Algorithm: Use Algorithm 1 to iteratively build the subset $\mathcal{A}$ by selecting data points that offer the highest marginal gain according to the selected submodular function.
By synergizing our novel utility-based kernel with submodular optimization, DELIFT achieves data-efficient fine-tuning that effectively addresses both redundancy and informativeness in the data selection process, optimizing the model’s performance across various tasks and domains.
4 Experimental Results
We conducted extensive experiments to evaluate the effectiveness of DELIFT across various fine-tuning scenarios, model scales, and datasets. This section details our experimental setup, baselines, evaluation metrics, and results analysis.
Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 37.87 78.92 2.98 36.36 82.55 3.02 25.76 43.34 1.42 35.50 80.46 2.58 Random 39.00 80.66 3.12 44.45 85.46 3.12 33.05 72.73 2.92 44.70 83.75 2.95 SelectIT 43.08 84.50 3.18 45.14 85.88 3.21 36.11 76.31 3.18 49.68 85.84 3.20 LESS 42.08 83.24 3.26 45.16 84.95 3.28 47.10 85.94 3.23 48.68 85.86 3.24 DELIFT (SE) 47.43 84.40 3.28 48.22 86.50 3.28 46.62 85.28 3.24 45.64 83.70 3.27 DELIFT 48.46 85.77 3.35 52.79 88.04 3.37 49.83 85.27 3.32 50.31 84.40 3.33 Full Data 58.65 88.72 3.45 65.51 92.24 3.51 55.92 88.26 3.45 74.98 93.33 3.84
Table 1: Results on Use Case 1: MixInstruct. Bold indicates the best performance. There is a 10.44% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.27% performance percentage drop from DELIFT to the next best baseline.
Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 18.03 59.13 1.54 20.15 58.38 1.78 20.10 48.66 1.36 20.64 49.17 1.39 Random 20.05 59.39 1.79 20.29 59.39 1.83 20.83 49.92 2.24 24.51 53.41 2.36 SelectIT 31.38 71.08 2.86 32.96 74.76 2.90 35.37 66.67 2.52 38.98 69.84 2.54 LESS 34.59 83.23 3.07 35.03 83.37 3.50 39.69 72.12 3.17 40.32 70.89 3.24 DELIFT (SE) 34.69 83.31 3.43 35.46 83.43 3.53 37.07 71.49 3.52 38.13 79.68 3.74 DELIFT 35.48 83.69 3.58 35.60 83.64 3.54 40.66 84.00 3.68 41.91 84.53 3.76 Full Data 36.43 84.25 3.53 35.88 76.87 3.63 42.07 85.26 3.78 44.73 87.03 3.82
Table 2: Results on Use Case 1: P3. Bold indicates the best performance. There is only a 0.76% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.23% performance percentage drop from DELIFT to the next best baseline.
4.1 Datasets and Use Cases
We evaluated DELIFT across the three previously described fine-tuning scenarios:
Use Case 1: Instruction Tuning We evaluated the effectiveness of DELIFT for use case 1 on two datasets: MixInstruct (Jiang et al., 2023) and P3 (Public Pool of Prompts) (Sanh et al., 2021). We randomly selected 21,000 train, 6,000 valid, and 3,000 test samples. Using the Facility Location (FL) submodular function, we aimed to select a subset of training data that was both representative and informative.
Use Case 2: Task-Specific Fine-Tuning We evaluated DELIFT for task-specific fine-tuning using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct with MT-Bench (Zheng et al., 2023). We used the Facility Location Mutual Information (FLMI) submodular function to select the most informative samples from the training datasets (HotpotQA and MixInstruct) that shared relevant information with the target datasets (MMLU and MT-Bench, respectively).
Use Case 3: Continual Fine-Tuning We evaluated DELIFT in a continual fine-tuning setting using two dataset pairs: (1) SQuAD (Rajpurkar et al., 2016) paired with HotpotQA for general question-answering, and (2) proprietary query rewriting datasets covering IBM and government domains. [^1] In this task, non-standalone questions –questions that require previous context to answer– must be rewritten to be standalone. For example, ”How much is it?” should be rewritten to ”How much is the subscription for IBM Cloud?” Such queries are common in user-agent conversations where a user asks a follow-up to an agent. Our goal was to integrate new knowledge efficiently while minimizing redundancy. We employed the Facility Location Conditional Gain (FLCG) submodular function, selecting complementary samples from the new dataset (HotpotQA and Government query rewrite) that provided additional, non-overlapping information to the existing dataset (SQuAD and IBM query rewrite).
4.2 Experimental Setup
Models: We evaluated DELIFT on two state-of-the-art open-source models: Phi-3-mini-128k-instruct (Abdin et al., 2024): 3.8B parameters, Qwen2-72B-Instruct (Yang et al., 2024): 72B parameters. These models were chosen to demonstrate effectiveness across different model scales.
Model Qwen2 Phi-3 Method QLoRA QLoRA Initial 82.10 69.10 Random 79.31 65.16 SelectIT 79.13 65.24 LESS 80.35 66.72 DELIFT (SE) 80.10 66.36 DELIFT 81.70 68.70 Full Data 78.36 64.50
Table 3: Results on Use Case 2: HotpotQA and MMLU (5-shot) for Qwen2 and Phi-3 models (classification accuracy). Bold indicates the best performance. For Qwen2, DELIFT outperforms Full Data by 3.34%, while for Phi-3, it improves by 4.20%.
Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 44.32 74.86 2.48 47.65 77.92 2.72 39.57 69.43 2.31 42.89 72.76 2.53 Random 49.78 79.54 2.83 52.91 82.67 3.05 44.63 74.28 2.62 47.85 77.39 2.84 SelectIT 54.92 83.71 3.12 57.86 86.59 3.31 49.75 78.64 2.91 52.68 81.52 3.13 LESS 59.63 85.89 3.29 62.74 88.72 3.48 54.82 81.95 3.08 57.73 84.67 3.29 DELIFT (SE) 62.85 86.94 3.38 65.83 89.76 3.57 57.69 82.87 3.17 60.54 85.59 3.38 DELIFT 64.73 87.82 3.47 67.91 90.64 3.66 59.58 83.76 3.26 62.47 86.48 3.47 Full Data 65.89 88.65 3.55 69.72 91.53 3.74 60.76 84.59 3.34 64.31 87.42 3.55
Table 4: Results on Use Case 2: MixInstruct and MT-Bench. Bold indicates the best performance. There is a 2.91% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 1.14% performance percentage drop from DELIFT to the next best baseline.
Metrics: We use a variety of metrics to characterize performance. For $n$-gram word overlap we use ROUGE (Lin, 2004). For semantic similarity we calculate the dot product between the embeddings from the bge-large-en-v1.5 model (Xiao et al., 2023); the embeddings are normalized to unit vectors, hence the closer the dot product is to 1, the more semantically similar the vectors (the metric is referred to as ‘BGE’). Additionally, we use Prometheus (Kim et al., 2023), specifically the prometheus-7b-v2.0 model, as an LLM-as-a-Judge (referred to as ‘LAJ’). With our custom rubric outlined in Appendix B, Prometheus assigns scores in a range of 1 to 5 (higher scores indicate better performance.) Finally, we use classification accuracy to evaluate MMLU.
Baselines: We evaluated DELIFT by comparing it against several baselines to understand its effectiveness in data selection. These baselines included: (1) SelectIT (Liu et al., 2024), which selects data using model feedback at the token, sentence, and model levels to identify useful samples; (2) LESS (Xia et al., 2024), which leverages LoRA approximated gradient-based influence estimation to prioritize impactful data points; (3) Random, which selects a fixed percentage (x%) of the dataset randomly, providing a benchmark for non-strategic selection; (4) DELIFT with Sentence Embedding Features (SE), which uses DELIFT but substitutes sentence embeddings as the feature space, employing a model-independent, pairwise similarity kernel instead of the utility kernel for submodular optimization; and (5) Full Data, where the entire dataset is used for fine-tuning, serving as an upper benchmark for performance. For In-Context Learning (ICL), the selected subsets from each baseline were used as the pool of examples, allowing us to evaluate how effectively each method supports ICL by providing relevant and informative data.
Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 44.11 70.49 2.43 48.49 80.85 2.62 40.66 58.68 1.52 43.96 69.56 2.29 Random 55.57 85.26 2.91 55.52 85.53 2.94 45.76 76.19 2.45 58.94 82.41 2.89 SelectIT 63.07 86.38 3.18 65.42 87.50 3.20 63.49 85.27 2.96 64.09 85.07 3.16 LESS 64.28 85.41 3.29 69.85 89.33 3.45 66.01 87.20 3.19 67.53 88.17 3.22 DELIFT (SE) 61.07 85.16 3.45 74.05 92.47 3.58 68.84 88.46 3.32 69.30 88.62 3.35 DELIFT 69.49 87.94 3.60 74.19 92.23 3.65 74.11 89.41 3.57 74.38 91.55 3.57 Full Data 66.08 87.84 3.65 76.83 92.63 3.74 71.23 91.10 3.52 77.12 91.10 3.64
Table 5: Results on Use Case 3: IBM and Government. Bold indicates the best performance. There is only a 0.31% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.89% performance percentage drop from DELIFT to the next best baseline.
Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 51.51 66.97 1.77 54.18 78.27 2.50 40.42 58.23 1.26 40.94 58.12 1.29 Random 54.38 79.12 2.57 59.23 82.02 2.66 44.29 59.45 1.33 50.29 61.52 1.60 SelectIT 58.03 83.75 2.82 63.26 84.01 2.87 47.35 74.15 2.54 56.88 80.47 2.70 LESS 67.16 85.76 2.94 69.72 86.63 3.26 60.97 81.41 2.84 61.56 81.53 2.88 DELIFT (SE) 73.75 88.01 3.26 74.84 88.79 3.30 64.44 83.95 3.03 66.35 84.77 3.14 DELIFT 76.94 90.41 3.33 77.56 89.99 3.34 66.55 84.65 3.25 67.09 85.17 3.32 Full Data 77.78 90.31 3.35 78.72 90.77 3.48 68.47 85.93 3.33 70.48 86.06 3.44
Table 6: Results on Use Case 3: SQuAD and HotpotQA. Bold indicates the best performance. There is only a 1.94% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.78% performance percentage drop from DELIFT to the next best baseline.
4.3 Results and Analysis
To ensure a fair and comprehensive evaluation of DELIFT, we conducted experiments across three distinct fine-tuning scenarios: instruction tuning, task-specific fine-tuning, and continual fine-tuning. For all subset selection methods—including DELIFT, Random, SelectIT, LESS, and DELIFT with Sentence Embdedding Features (SE)—we consistently selected 30% of the dataset as a subset, enabling direct comparisons between methods and with the full dataset baseline (see Section 4.4 for an ablation study examining the impact of subset size).
Use Case 1: Instruction Tuning Our first set of experiments focused on instruction tuning, a crucial task to enhancing a model’s ability to follow diverse instructions. As shown in Tables 1 and 2, DELIFT achieved a minimal performance drop of only 5.60% compared to using the full dataset while reducing the dataset by 70%. This demonstrates DELIFT’s capability to retain the most informative samples essential for instruction tuning. Furthermore, DELIFT outperformed other subset selection methods, achieving a 2.74% improvement and a substantial 26.21% advantage over the next best and worst-performing baselines, respectively. These results underscore DELIFT’s superior ability to maintain high performance with significantly reduced data, highlighting its efficacy in instruction tuning.
Use Case 2: Task-Specific Fine-Tuning In the task-specific fine-tuning scenario, we evaluated DELIFT using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct paired with MT-Bench (Zheng et al., 2023). Results, presented in Tables 4 and 3, demonstrate DELIFT’s consistent and competitive performance across different task pairs. A particularly noteworthy outcome emerged from the HotpotQA-MMLU pair, where DELIFT not only matched but exceeded the performance of the full dataset, achieving a 5.51% improvement. This indicates that DELIFT’s selective approach can effectively filter out noise and focus on the most relevant and informative samples, yielding enhanced task-specific adaptation even with reduced data.
Use Case 3: Continual Fine-Tuning The third use case examined DELIFT’s efficacy in continual fine-tuning, where models need to incorporate new information while retaining previously learned knowledge. As detailed in Tables 5 and 6, DELIFT demonstrated remarkable consistency, showing only a marginal 1.13% performance drop compared to using the full dataset. Moreover, DELIFT outperformed the second-best baseline by 3.33% and the worst baseline by 23.88%, highlighting its superiority in data selection. In specialized tasks such as query rewriting, DELIFT even surpassed the performance of the full dataset, suggesting that its selective approach effectively prunes noisy or irrelevant data points, thereby enhancing model performance.
4.4 Ablation Study: Impact of Subset Size
To assess how subset size influences DELIFT’s performance, we conducted an ablation study by varying the subset size from 5% to 100% of the full dataset across three use cases. The results, detailed in Appendix A and illustrated in Figure 2, show that LAJ scores generally increase with subset size. Utilizing the full dataset consistently yields the highest performance, highlighting the benefits of larger training sets. However, for methods such as DELIFT, SelectIT, and LESS, performance gains plateau or slow beyond a 50% subset size, indicating that additional data offers minimal benefits and may introduce redundancy. Importantly, DELIFT outperforms all baselines across subset sizes from 5% to 100%, demonstrating its robustness and effectiveness in selecting informative samples regardless of subset size. These findings suggest that carefully selected smaller datasets can achieve comparable performance to larger, unfiltered datasets, which is particularly valuable for resource-intensive large language models.
4.5 Discussion
The comprehensive results across all three use cases highlight DELIFT’s effectiveness and versatility. By consistently reducing data requirements by up to 70% while maintaining—and in some cases improving—performance, DELIFT addresses a critical challenge in large language model fine-tuning. The superior performance of DELIFT can be attributed to its novel pairwise utility metric and the use of tailored submodular functions for each fine-tuning stage. This approach enables DELIFT to select not only representative and diverse samples but also to reduce noise present in the full dataset. The ability to outperform full datasets in certain scenarios, particularly in niche tasks like query rewriting, underscores DELIFT’s capacity to distill the most relevant and informative data points. These findings have significant implications for the accessibility and efficiency of LLM fine-tuning. By dramatically reducing the amount of data required for effective fine-tuning, DELIFT paves the way for more widespread adoption and application of large language models across various domains, especially in resource-constrained environments. Furthermore, DELIFT’s consistent outperformance of existing data selection techniques across various fine-tuning scenarios and model scales demonstrates its robustness and broad applicability, making it a valuable tool for researchers and practitioners alike. In conclusion, our experimental results firmly establish DELIFT as a powerful and efficient method for data selection in LLM fine-tuning. By addressing the critical challenge of optimal data selection, DELIFT not only enhances the efficiency of model training but also opens new possibilities for fine-tuning large language models in domains where data or computational resources may be limited.
5 Conclusion, Limitations, and Future Work
In this paper, we introduced DELIFT, a novel approach to data-efficient fine-tuning of large language models by employing a versatile pairwise utility metric combined with submodular optimization techniques for optimal data selection. Empirical evaluations showed that DELIFT can reduce data and computational requirements by up to 70% while achieving performance comparable to the full dataset, and outperforming existing data selection methods by up to 26% in effectiveness. These results suggest that DELIFT offers a promising method for improving the accessibility of LLM adaptation, especially for resource-constrained scenarios. However, our approach has limitations, including potential sensitivity to the quality and diversity of initial data and the risk of bias amplification inherent in the selected data. Future work will explore integrating DELIFT with data augmentation techniques to improve robustness, incorporating fairness constraints to mitigate biases, and extending the approach to emerging model architectures and multimodal learning. Our ongoing efforts are directed toward ensuring that DELIFT contributes to responsible and equitable AI development while maximizing efficiency.
6 Acknowledgement
This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS240550 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Appendix
Appendix A Subset Size Comparison
To assess how subset size influences the performance of DELIFT, we performed an ablation study by varying the subset size from 5% to 100% (specifically 5%, 15%, 30%, 50%, 100%) of the entire dataset across three distinct use cases. Figure 2 illustrates the performance metric LAJ as a function of subset size for each fine-tuning scenario.
A.1 General Observations
  • Performance Increases with Subset Size: Across all methods, LAJ scores generally improve as the subset size increases. Utilizing the full dataset consistently yields the highest performance, underscoring the benefits of a larger training set.
  • Diminishing Returns Beyond 50%: For methods such as DELIFT, SelectIT, and LESS, performance gains plateau or slow down beyond a 50% subset size. This suggests that additional data beyond this point offers minimal benefits and may introduce redundancy.
A.2 Detailed Analysis of Methods
A.2.1 Initial vs. Random Selection
  • Initial Baseline: Consistently records the lowest scores across all subset sizes, indicating that models without data-informed selection struggle to generate quality responses.
  • Random Selection: Slightly outperforms the Initial baseline but maintains a relatively flat performance curve. This lack of significant improvement highlights that uninformed data selection does not substantially enhance model quality.
A.2.2 SelectIT and LESS Methods
  • LESS: Demonstrates a strong upward trend, particularly when subset sizes increase from 15% to 50%. This indicates that LESS effectively selects informative subsets, especially in the mid-range subset sizes, but is sub-optimal with smaller subset sizes.
  • SelectIT: Initially lags behind DELIFT and LESS but shows steady improvement with larger subset sizes. For subset sizes above 50%, SelectIT approaches the performance of DELIFT, suggesting its heuristic-driven selection becomes more effective with more data.
A.2.3 DELIFT Variants
  • DELIFT vs. DELIFT (SE): DELIFT consistently outperforms DELIFT (SE), which uses sentence embeddings, highlighting the superiority of DELIFT’s utility-based kernel in capturing data informativeness.
  • DELIFT vs. Other Methods: DELIFT outperforms all other subset selection methods across all subset sizes, particularly below 50%. This effectiveness is attributed to DELIFT’s strategy of identifying the most informative samples early on, making it ideal for scenarios with limited computational resources.
  • DELIFT vs. Full Data: At smaller subset sizes (e.g., 15%, 30%), DELIFT achieves LAJ scores close to the Full Data baseline. In ICL fine-tuning scenarios, a 30% subset size with DELIFT nearly matches Full Data performance, demonstrating its efficiency in data reduction without significant loss in performance.
A.3 Impact on Different Fine-Tuning Scenarios
  • ICL vs. QLoRA: QLoRA fine-tuning generally yields higher scores than ICL across all methods, suggesting that QLoRA benefits more from effective data selection strategies. DELIFT, in particular, shows more pronounced improvements in QLoRA settings, indicating its subsets are well-suited for efficient parameter tuning.
  • Use Case Comparisons: In Use Case 3 (IBM and Government datasets), DELIFT achieves the highest gains relative to the Initial baseline across both ICL and QLoRA scenarios. This effectiveness is likely due to the nature of query rewriting tasks, where DELIFT’s informed data selection effectively eliminates redundant or irrelevant examples, resulting in a higher-quality training set.
Figure 2: Graphs of LLM-A-J scores (y-axis) of Qwen2-72B-Instruct with varying subset sizes (x-axis) of Use Case 1 on MixInstruct for (a) ICL and (b) QLoRA, Use Case 2 on MixInstruct and MT-Bench for (c) ICL and (d) QLoRA, and Use Case 3 on IBM and Government for (e) ICL and (f) QLoRA.
a)Refer to caption b)Refer to caption c)Refer to caption d)Refer to caption e)Refer to caption f)Refer to caption
Appendix B Prometheus Rubric
The Prometheus model served as an LLM-as-a-Judge to evaluate response quality from different data selection methods. Table 7 contains the general rubric used for the Prometheus model scoring on all use cases and settings (except for the experiments on the query-rewriting task using the IBM-proprietary data).
Evaluate the model’s ability to follow instructions and deliver a high-quality response across the following dimensions:
1. Instruction Following: How accurately and fully does the model adhere to the given instruction?
2. Accuracy: Is the information correct, reliable, and factually sound?
3. Relevance: Does the response directly address the question or task without unnecessary information?
4. Completeness: Does the response cover all essential aspects of the instruction or question
5. Depth: How thoroughly does the response explore the topic? Does it demonstrate insightful analysis where appropriate?
6. Clarity: Is the response well-organized, easy to follow, and free from ambiguity or confusion?
7. Creativity: Does the response offer original or innovative approaches where applicable?
8. Helpfulness: Does the response effectively meet the user’s needs and provide value in solving the problem or addressing the query?
Score of 1: The response fails to meet expectations across most or all criteria. It does not follow the instruction, contains significant errors or misinformation, lacks relevance, is incomplete or shallow, unclear, unoriginal, and unhelpful.
Score of 2: ”The response shows major deficiencies across several criteria. It partially follows the instruction but includes significant inaccuracies, is often irrelevant, incomplete, or lacks depth, clarity, creativity, and helpfulness.
Score of 3: ”The response is average, meeting some but not all criteria. It follows the instruction but may fall short in terms of accuracy, depth, relevance, or helpfulness. Improvements in clarity and insightfulness may be needed.
Score of 4: The response is strong, performing well across most criteria. It follows the instruction closely, is mostly accurate and relevant, provides good depth, and is well-structured. Minor improvements could enhance clarity, creativity, or helpfulness.
Score of 5: ”The response excels in all or nearly all criteria. It fully follows the instruction, is highly accurate, directly relevant, complete, and demonstrates depth and insight. The response is well-organized, creative where appropriate, and very helpful in addressing the user’s needs.
Table 7: General Prometheus Rubric
B.1 Usage Notes
  • Each response is evaluated independently based on the criteria above.
  • The cumulative score reflects the overall quality and effectiveness of the response.
  • Final LAJ scores are obtained by averaging the scores across all criteria.
Appendix C LLM-as-Judges Scores
In Tables 8 and 9, we show the distribution of Prometheus scores on one particular setting: Use Case 1, MixInstruct training and MixInstruct validation sets on the Qwen2-72B-Instruct model. These figures make clear that the average LGA scores computed in Tables 1-6 are true averages of a distribution of scores, not averages of a combination of just 1’s and 5’s.
ICL QLoRA Initial [Uncaptioned image] [Uncaptioned image] Random [Uncaptioned image] [Uncaptioned image] SelectIT [Uncaptioned image] [Uncaptioned image]
Table 8: LLM-as-Judges score distributions for Use Case 1 with MixInstruct training and validation set on the Qwen2-72B-Instruct model on the Initial, Random, and SelectIT baselines. The corresponding table is Table 1.
ICL QLoRA LESS [Uncaptioned image] [Uncaptioned image] DELIFT (SE) [Uncaptioned image] [Uncaptioned image] DELIFT [Uncaptioned image] [Uncaptioned image] Full Data [Uncaptioned image] [Uncaptioned image]
Table 9: LLM-as-Judges score distributions for Use Case 1 with MixInstruct training and validation set on the Qwen2-72B-Instruct model on the LESS, DELIFT with Sentence Embedding, DELIFT, and Full Data methods. The corresponding table is Table 1.
C.1 Interpretation of Score Distributions
C.1.1 Overall Trends
  • Score Variability: There is significant variability in score distributions across different methods. The Initial and Random baselines show a concentration of scores between 2.5 and 3.5, indicating average to subpar performance.
  • Enhanced Performance with Advanced Methods: Methods like SelectIT, LESS, DELIFT (SE), and DELIFT exhibit score distributions skewed towards higher values (3.5 to 4.0), with DELIFT showing the highest concentration above 3.5. This highlights their effectiveness in selecting informative and useful data for fine-tuning.
C.1.2 Method-Specific Observations
  • Initial and Random Methods: Both methods have lower mean scores (around 3.0 to 3.2) with wide spreads, suggesting inconsistent and generally lower-quality responses.
  • SelectIT and LESS Methods:
    • SelectIT: Shows improved mean scores, especially in QLoRA settings, indicating its effectiveness in resource-constrained training scenarios.
    • LESS: Demonstrates significant performance improvements, with mean scores around 3.26 to 3.28, reflecting effective gradient-based data selection.
  • DELIFT Variants:
    • DELIFT (SE): Skews towards higher scores but not as prominently as DELIFT.
    • DELIFT: Achieves the highest average scores (3.35 for ICL and 3.37 for QLoRA), outperforming all other methods and indicating its superior utility-based kernel and submodular optimization.
C.1.3 Comparison with Full Data
  • DELIFT vs. Full Data: DELIFT nearly matches Full Data performance with only a slight reduction in mean scores (3.35 to 3.37 vs. 3.45 to 3.51). This demonstrates DELIFT’s capability to retain most of the model’s performance while using significantly less data.
  • Efficiency of Data Pruning: Full Data shows a modest increase in mean scores compared to DELIFT, but at the cost of substantially higher computational resources. DELIFT offers a more efficient alternative without major sacrifices in performance.
Appendix D Limitations
  • Dependence on Initial Data Quality: DELIFT’s effectiveness relies on the diversity and quality of the initial dataset. Biases or lack of diversity in the dataset can propagate to the selected subsets.
  • Scalability Constraints: While DELIFT is computationally efficient, extremely large datasets may still present challenges in terms of computation and memory.
  • Domain-Specific Performance: DELIFT’s performance may vary across different domains, particularly those requiring specialized knowledge or handling multimodal data.
  • Bias Amplification Risks: The subset selection process may unintentionally amplify existing biases within the data, necessitating careful mitigation strategies.
Appendix E Future Work
  • Integration with Data Augmentation: Combining DELIFT with data augmentation techniques could further enhance the robustness and diversity of selected subsets.
  • Fairness and Bias Mitigation: Incorporating fairness constraints and bias mitigation strategies into the subset selection process to ensure equitable model performance across different groups.
  • Extension to Multimodal Learning: Adapting DELIFT for multimodal data (e.g., text, images, audio) to expand its applicability beyond natural language processing.
  • Theoretical Analysis: Developing a deeper theoretical understanding of the utility metric and its properties to further validate and refine the approach.
  • Enhancing Scalability: Exploring methods to scale DELIFT effectively for larger datasets and more complex models without compromising efficiency.
Our ongoing efforts aim to ensure that DELIFT contributes to responsible and equitable AI development while maximizing efficiency.
Appendix F Code and Data Availability
To facilitate reproducibility and further research, we will make the DELIFT implementation and the datasets used in our experiments publicly available upon publication. Interested researchers can access these resources through the following repository: https://anonymous.4open.science/r/optimizing-data-selection-0CD0.
Appendix G Hyperparameter Settings
Consistent hyperparameter settings were maintained across all experiments to ensure reproducibility:
  • Submodular Function: Utilized Facility Location (FL), Facility Location Mutual Information (FLMI), or Facility Location Conditional Gain (FLCG) based on the use case.
  • Utility Metric Scaling Factor: Set $\eta=1$ for FLMI and $\nu=1$ for FLCG.
  • Budget (of Data): Fixed at 30% for all subset selection experiments.
  • Optimization Algorithm: Employed greedy maximization with a stopping criterion based on the budget.
  • Distance Metric: Used length-normalized L2 norm.
  • Teacher Forcing Technique: Applied during utility metric computation to ensure reliable prediction accuracy measurement.

### v3

@@
 Abstract
-Fine-tuning large language models (LLMs) is essential for enhancing their performance on specific tasks but is often resource-intensive due to redundant or uninformative data.
-To address this inefficiency, we introduce DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel algorithm that systematically optimizes data selection across the three key stages of fine-tuning: (1) instruction tuning, (2) task-specific fine-tuning (e.g., reasoning, question-answering), and (3) continual fine-tuning (e.g., incorporating new data versions).
-Unlike existing methods that focus on single-stage optimization or rely on computationally intensive gradient calculations, DELIFT operates efficiently across all stages.
-Central to our approach is a pairwise utility metric that quantifies how beneficial a data sample is for improving the model’s responses to other samples, effectively measuring the informational value relative to the model’s current capabilities.
-By leveraging different submodular functions applied to this metric, DELIFT selects diverse and optimal subsets that are useful across all stages of fine-tuning.
-Experiments across various tasks and model scales demonstrate that DELIFT can reduce the fine-tuning data size by up to 70% without compromising performance, offering significant computational savings and outperforming existing methods in both efficiency and efficacy.
+Fine-tuning large language models (LLMs) is crucial for task specialization but often becomes resource-intensive due to redundant or uninformative data.
+Existing data selection methods typically rely either on computationally expensive gradient-based metrics or static embeddings that fail to adapt dynamically to the model’s evolving state, thus limiting their practical effectiveness.
+To address this, we propose DELIFT (Data Efficient Language model Instruction Fine-Tuning), leveraging a novel, computationally efficient utility metric inspired by In-Context Learning (ICL).
+Our ICL-based metric measures the informational value of each data sample by quantifying its effectiveness as an in-context example in improving model predictions for other samples, reflecting its actual contribution relative to the model’s current state.
+Integrated with tailored submodular optimization methods, DELIFT systematically selects diverse, informative subsets optimized specifically for each fine-tuning stage: instruction tuning, task-specific adaptation, and continual fine-tuning.
+Experimental results across multiple datasets and model scales show DELIFT reduces fine-tuning data requirements by up to 70% without compromising performance, consistently outperforming existing methods by up to 26% in effectiveness and efficiency.
 1 Introduction
-Fine-tuning large language models (LLMs) is pivotal for adapting these powerful architectures (Devlin et al., 2019; Brown et al., 2020a; Touvron et al., 2023) to specialized tasks such as intricate reasoning, precise question-answering, and the seamless integration of new information (Ouyang et al., 2022).
-This transformation—from a general-purpose model to a task-specific agent—heavily relies on the quality and nature of the data employed during fine-tuning, which critically determines the model’s subsequent performance (Wei et al., 2022; Zhou et al., 2023; Hoffmann et al., 2024).
-The effectiveness of fine-tuning hinges on the quality, diversity, and relevance of the selected data (Gururangan et al., 2020; Wei et al., 2022; Zhou et al., 2023).
-High-quality data ensures accurate learning, diverse data enhances generalization, and relevant data aligns the model’s capabilities with specific application needs.
-However, optimizing data selection across different fine-tuning phases remains a significant challenge, leading to our central research question:
-How can we create a unified framework for efficient data selection across all fine-tuning stages of LLMs, while optimizing performance and maximizing data efficiency?
-To address this challenge, we present DELIFT (Data Efficient Language model Instruction Fine-Tuning), a novel, unified, and computationally efficient algorithm engineered to optimize data selection across all stages of the fine-tuning process.
-The key innovation of DELIFT lies in its pairwise utility metric, which assesses the informational value of data samples relative to both the model’s current capabilities and other samples within the dataset.
-This metric, combined with submodular optimization techniques, allows DELIFT to efficiently select optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs.
-The typical fine-tuning process comprises three key stages: 1.
-Instruction Tuning: Enhances the model’s ability to follow general instructions (Mishra et al., 2022; Wei et al., 2022; Longpre et al., 2023); 2.
-Task-Specific Fine-Tuning: Refines the model’s expertise in specific domains (Gururangan et al., 2020; Cobbe et al., 2021); 3.
-Continual Fine-tuning: Enables the model to integrate new information while mitigating catastrophic forgetting (Madotto et al., 2021; Wu et al., 2024).
-DELIFT is able to optimize data selection processes across all three stages.
-Additionally, DELIFT offers significant benefits for In-Context Learning (ICL) (Brown et al., 2020b; Xue et al., 2024).
-By utilizing the selected subsets as the ICL example pool, DELIFT achieves similar or better performance compared to using the entire dataset, thereby enhancing data efficiency in ICL scenarios.
-This dual functionality is empirically validated in our experimental results.
-Existing data selection methodologies often fail to address the nuanced requirements of the aforementioned distinct fine-tuning stages.
-Many approaches are tailored to a single stage, lacking the adaptability needed for comprehensive fine-tuning (Xia et al., 2024; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024).
-Others depend on computationally intensive procedures, such as exhaustive gradient computations, rendering them impractical for large-scale models and datasets (Killamsetty et al., 2021b; a; Xia et al., 2024; Zhang et al., 2024).
-Additionally, some methods utilize features obtained from an independent model that are not specifically aligned with the model undergoing fine-tuning, reducing their effectiveness (Killamsetty et al., 2023; Liu et al., 2024; Bukharin & Zhao, 2024; Chen et al., 2024; Du et al., 2023).
-DELIFT addresses these limitations by adapting to the unique requirements of each fine-tuning stage.
-1.
-Instruction Tuning: Selects diverse data to enhance general instruction-following capabilities; 2.
-Task-Specific Fine-Tuning: Prioritizes data that is aligned with the target task, to refine specialized expertise; 3.
-Continual Fine-tuning: Identifies novel, complementary information to expand the model’s knowledge base while safeguarding against catastrophic forgetting.
-Figure 1 illustrates how DELIFT optimizes data selection across these stages, demonstrating the selection and pruning processes in each fine-tuning phase.
-By leveraging submodular optimization techniques (Fujishige, 2005; Bilmes, 2022) and submodular information measures (Iyer et al., 2021), DELIFT efficiently selects optimal data subsets that precisely address the model’s learning requirements without incurring unnecessary computational costs.
-This approach effectively balances data utility and computational efficiency.
+Large Language Models (LLMs) have become indispensable for solving a variety of natural language processing tasks, ranging from question answering and summarization to complex dialogue and reasoning (Brown et al., 2020; Touvron et al., 2023).
+Despite their remarkable adaptability, fine-tuning LLMs often requires enormous computational resources and time, especially when a significant portion of the training data is either redundant or uninformative (Gururangan et al., 2020; Sorscher et al., 2023).
+This challenge grows more critical with increasing model and dataset sizes, posing a key limitation to the broader deployment of LLMs.
+Existing data selection methods generally fall under two paradigms: (1) static embedding-based approaches that compute sample similarities without reflecting the model’s evolving state (Bukharin & Zhao, 2024; Chen et al., 2024), and (2) gradient-based methods that offer more model-specific feedback but often entail prohibitive computational overhead, especially for large-scale models (Killamsetty et al., 2021b; Xia et al., 2024).
+Although both paradigms can yield initial benefits, they often fail to account for how a model’s knowledge shifts over multiple fine-tuning phases: (1) Instruction Tuning (Mishra et al., 2022; Wei et al., 2022; Longpre et al., 2023), which enhances the model’s ability to follow diverse instructions; (2) Task-Specific Fine-Tuning (Gururangan et al., 2020; Cobbe et al., 2021), which focuses on refining domain expertise; and (3) Continual Fine-Tuning (Madotto et al., 2021; Wu et al., 2024), which incrementally incorporates new knowledge while mitigating catastrophic forgetting.
+Thus, a natural question arises:
+Can we develop a unified, computationally efficient data selection framework that adapts to all stages of fine-tuning and maximizes model performance while minimizing data redundancy?
 Figure 1: DELIFT data selection across fine-tuning stages.
@@
 a)Refer to caption b)Refer to caption c)Refer to caption
-Our key contributions are as follows:
-1) Versatile Pairwise Utility Metric: A novel, easy-to-compute metric for assessing data informativeness, incorporating model feedback applicable across all fine-tuning stages.
-2) Unified Data Selection Algorithm: DELIFT systematically optimizes data selection for instruction tuning, task-specific fine-tuning, and continual fine-tuning within a single framework.
-3) Computational Efficiency: Circumvents resource-intensive operations, ensuring scalability to large datasets and models.
-DELIFT achieves at least 70% reduction in computational time compared to gradient-based methods on benchmark tasks.
-4) Enhanced Performance with Reduced Data: Demonstrates the ability to reduce fine-tuning data size by up to 70% without compromising performance, and achieves comparable efficacy as to utilizing the full dataset.
-5) Improvement over Existing Methods: Outperforms current data selection techniques by up to 26% in effectiveness across diverse tasks and model scales (see Section  4).
-The remainder of this paper is organized as follows: Section 2 provides background on fine-tuning LLMs and reviews related work.
-Section 3 details the methodology behind DELIFT, including the development of our pairwise utility metric and the submodular optimization process.
-Section 4 presents experimental results that showcase the effectiveness and efficiency of our method.
-Section 5 discusses the implications of our findings and potential future directions.
-Finally, we release our code base for further research.
+In this paper, we introduce DELIFT (Data-Efficient Language Model Instruction Fine-Tuning), a single-stop solution designed to address data selection across all fine-tuning stages within a single framework.
+DELIFT is grounded in information theory yet uses the practical intuition of in-context examples to assess the ’information gain’ of each data sample relative to the current state of a model.
+Specifically, we propose a new utility metric that captures how effectively one sample improves the model’s prediction of another.
+By combining these pairwise utilities with submodular optimization, DELIFT generates diverse, nonredundant subsets uniquely tailored to each fine-tuning phase.
+We evaluated DELIFT on various tasks and model scales, consistently observing that it can prune up to 70% of the training data without hurting performance - and often improving it - outperforming existing methods by up to 26% in efficiency and effectiveness.
+In doing so, we show that careful, utility-driven data selection can be far more effective than sheer data volume, opening the door to more resource-friendly and targeted fine-tuning.
+Our primary contributions are as follows.
+1.
+A unified information-theoretic data selection paradigm that leverages pairwise utilities grounded in conditional pointwise mutual information, making it adaptable to instruction tuning, task-specific adaptation, and continual fine-tuning.
+2. A single-stop, submodular optimization framework that integrates these utilities to provide diverse, high-value subsets for each fine-tuning stage without incurring prohibitive computation.
+3.
+Extensive empirical validation showing up to 70% data reduction with minimal (and sometimes zero) performance loss across multiple domains, demonstrating substantial gains in both efficacy and efficiency.
+The remainder of this paper is organized as follows.
+Section 2 reviews prior work on data-efficient strategies for fine-tuning LLMs and situates our approach within the literature.
+Section 3 introduces our information-theoretic utility metric and describes how it integrates with submodular optimization to enable data selection across diverse fine-tuning stages.
+Section 4 presents comprehensive experiments demonstrating the effectiveness and efficiency of our framework on multiple tasks and models.
+Finally, Section 5 discusses the broader implications of our results, outlines limitations, and suggests directions for future research.
+Our complete code base is publicly available at https://github.com/agarwalishika/delift, enabling further exploration and replication.
 2 Related Work
-Efficient data subset selection is vital for enhancing training efficiency in deep neural networks while maintaining or improving model performance.
-This section categorizes existing subset selection methods into model-independent and model-dependent approaches and identifies the gaps our work addresses.
-Model-independent subset selection methods focus on selecting representative subsets without model-specific feedback.
-Common approaches include using pre-trained sentence embeddings with distance or clustering metrics (Bukharin & Zhao, 2024; Sorscher et al., 2023; Killamsetty et al., 2023; Du et al., 2023; Bhatt et al., 2024), as well as employing large models like GPT-4 or pre-trained reward models for high-quality data filtering (Du et al., 2023; Chen et al., 2024).
-However, these methods often struggle to translate the assessed diversity or quality into downstream utility.
-Model-dependent subset selection aims to identify data samples beneficial to the downstream model by analyzing features like per-sample gradients or loss values.
-Methods such as GradMatch (Killamsetty et al., 2021a), CRAIG (Mirzasoleiman et al., 2020), and TAGCOS (Zhang et al., 2024) focus on selecting samples that approximate the gradient updates over the full dataset.
-GLISTER (Killamsetty et al., 2021b) employs bilevel optimization to align gradients from selected subsets with those of a validation set.
-LESS (Xia et al., 2024) proposes computing gradients through LoRA fine-tuning to reduce the computational cost of gradient computation and utilizes random projection to address gradient dimensionality issues.
-Li et al.
-(2024) proposed the IFD score, a computationally efficient model-dependent metric that assesses instruction difficulty to filter challenging samples, though it does not guarantee data diversity.
-While effective in capturing useful samples, these methods often face computational challenges, especially with LLMs.
-Persistent limitations across these methods include: (i) Limited Adaptability across different fine-tuning stages, (ii) Computational Intensity due to model feedback reliance, (iii) Lack of Unified Solutions applicable across all fine-tuning phases, and (iv) ineffective Redundancy Handling.
-DELIFT addresses these limitations through a novel pairwise utility metric, which effectively aligns with the model’s evolving capabilities throughout fine-tuning.
-By integrating submodular optimization with pairwise model-dependent metrics that evaluate relative sample utility, DELIFT minimizes redundancy while maximizing adaptability and computational efficiency.
-This approach proves effective across diverse use cases including instruction tuning, task-specific fine-tuning, continual fine-tuning, and In-Context Learning (ICL), offering a versatile and scalable solution for data subset selection.
+Data Subset Selection for Deep Neural Networks.
+Selecting an informative subset of training samples is a longstanding strategy to reduce computational costs and enhance model generalization.
+Model-Independent Approaches.
+Traditional model-independent techniques, such as clustering or distance metrics on pre-trained embeddings (Bukharin & Zhao, 2024; Du et al., 2023; Killamsetty et al., 2023), capture broad semantic similarities but do not reflect the model’s changing state, limiting their effectiveness during iterative fine-tuning.
+Model-Dependent Approaches.
+Model-dependent methods incorporate the model’s evolving knowledge by analyzing gradients or loss values (Killamsetty et al., 2021b; a; Xia et al., 2024), often outperforming static approaches.
+However, performing gradient or influence estimations at scale becomes prohibitively expensive for large models.
+Techniques like LESS (Xia et al., 2024) alleviate some overhead via parameter-efficient fine-tuning (e.g., LoRA), , yet still incur repeated gradient or influence calculations that scale poorly with dataset size.
+Subset Selection with LLM Feedback.
+Another emerging direction leverages LLM feedback to score or filter training samples.
+For instance, SelectIT (Liu et al., 2024) employs self-reflection prompts to rate data quality, while filtering approaches using GPT-4 (Chen et al., 2024) rely on external heuristics.
+Though these provide a form of model-aware sampling, they typically lack a principled theoretical grounding.
+In addition, all these approaches primarily target a single fine-tuning stage, limiting their adaptability for instruction tuning, task-specific adaptation, or continual learning.
+Our Contribution.
+In contrast, we present a unified, information-theoretic framework that operates effectively across all fine-tuning stages: instruction tuning, task-specific adaptation, and continual fine-tuning.
+Our novel utility metric quantifies how one data point aids the prediction of another, mirroring the model’s evolving knowledge.
+Integrated within a submodular selection paradigm (Fujishige, 2005; Iyer et al., 2021), this approach balances diversity, coverage, and informativeness throughout the entire fine-tuning pipeline.
+As a result, we bridge the gap left by existing methods that are either restricted to a single phase or computationally infeasible at scale, demonstrating consistent performance improvements and notable efficiency gains.
 3 Methodology
-This section presents foundational concepts and the specific approach of DELIFT, focusing on data subset selection through a utility-based kernel integrated with submodular optimization techniques.
-3.1 Notation
-Let $\mathcal{D}$ denote the fine-tuning dataset, comprising elements $d_{i}=(x_{i},y_{i})$, where $x_{i}$ is the input sequence and $y_{i}$ is the corresponding output sequence.
-Our objective is to select a subset $\mathcal{A}\subseteq\mathcal{D}$ that maximizes the model’s performance while minimizing computational costs.
-The selection strategy adapts based on the fine-tuning objective, which may include instruction tuning, task-specific adaptation, or continual learning.
-3.2 Utility-Based Kernel
-At the core of DELIFT lies the utility-based kernel, a mechanism designed to quantify the informativeness of one data point when used as an in-context example for another.
-Consider two data points, $(x_{i},y_{i})$ and $(x_{j},y_{j})$.
-The utility of data point $j$ relative to data point $i$, denoted as $UF_{ij}$, is defined as:
-$UF_{ij}=d(GT_{i},p(y_{i}\mid x_{i}))-d(GT_{i},p(y_{i}\mid x_{j},y_{j},x_{i})),$ (1)
-where: $d(\cdot,\cdot)$ is a length-normalized distance metric between two probability distributions, $GT_{i}$ is the ground truth distribution for the sequence $y_{i}$, modeled as a vector of ones for each token to signify perfect prediction, $p(y_{i}\mid x_{i})$ is the model’s predicted probability distribution for $y_{i}$ given only the input $x_{i}$, $p(y_{i}\mid x_{j},y_{j},x_{i})$ is the predicted distribution for $y_{i}$ when the model is provided with $(x_{j},y_{j})$ as an in-context example followed by $x_{i}$.
-The distance metric $d(p,q)$ is calculated using the length normalized L2 norm and is defined as:
-$d(p,q)=\sqrt{\frac{\sum_{k=1}^{N}(p_{k}-q_{k})^{2}}{N}},$ (2)
-where $p_{k}$ and $q_{k}$ are the $k$-th elements of the flattened probability distributions $p$ and $q$, respectively.
-Importantly, $N$ varies with each data sample and corresponds to the number of tokens in the ground truth response $y_{i}$.
-This normalization ensures that the distance measure remains scale-invariant across different sequence lengths and vocabulary sizes.
-To compute the probability distributions accurately, we employ the teacher forcing technique (Williams & Zipser, 1989).
-This method ensures that the model uses the ground truth of previous tokens when predicting each subsequent token in the sequence, enabling reliable measurement of prediction accuracy.
-The utility value $UF_{ij}$ measures the improvement in prediction accuracy for the entire sequence $(x_{i},y_{i})$ when utilizing $(x_{j},y_{j})$ as an in-context example.
-A positive $UF_{ij}$ indicates that including data point $j$ enhances the model’s prediction accuracy for $i$, whereas a negative value suggests an adverse effect.
-3.3 Submodular Functions for Dataset Selection
-To optimize the selection of informative data subsets, DELIFT leverages submodular functions (Fujishige, 2005).
-Submodular functions are characterized by the property of diminishing marginal returns, making them ideal for selecting diverse, informative, and non-redundant subsets.
-Submodular function maximization can be efficiently approximated using a greedy algorithm, with a provable approximation guarantee of $1-\frac{1}{e}$ of the optimal solution (Nemhauser et al., 1978).
-We employ three tailored submodular functions (Iyer et al., 2021), each suited to a specific fine-tuning stage:
-3.3.1 Facility Location (FL)
-From an information perspective, the Facility Location function maximizes the coverage of the information space.
-It ensures that the selected subset $\mathcal{A}$ contains examples that are collectively representative of the entire dataset’s $\mathcal{D}$ information content.
-This is particularly useful in instruction tuning, where we aim to capture a diverse range of instruction types and their informational content.
-It is defined as the following where $s_{ij}$ is the similarity measure between data points $i$ and $j$:
-$f_{FL}(\mathcal{A})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij},$ (3)
-3.3.2 Facility Location Mutual Information (FLMI)
-The FLMI function is designed to maximize the mutual information between the selected subset $\mathcal{A}$ and the target domain dataset $\mathcal{D_{T}}$.
-In our context, it ensures that the selected data points are not only informative in general but also particularly relevant to the specific task at hand.
-This makes it ideal for task-specific fine-tuning, where we want to bridge the gap between general knowledge and task-specific information.
-It is defined below where $\eta$ is a scaling factor (set to $1$ in our experiments):
-$f_{FLMI}(\mathcal{A};\mathcal{D}_{T})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}+\eta\sum_{j\in\mathcal{A}}\max_{i\in\mathcal{D}_{T}}s_{ij},$ (4)
-3.3.3 Facility Location Conditional Gain (FLCG)
-From an information-theoretic standpoint, the FLCG function aims to maximize the conditional information gain of the selected subset $\mathcal{A}$ given the existing dataset $\mathcal{D_{E}}$.
-It quantifies how much new information each data point brings, conditional on what the model already knows.
-This is crucial for continual fine-tuning, where we want to avoid redundancy and focus on novel, complementary information that expands the model’s knowledge base without unnecessary repetition.
-It is defined as the following where $\nu$ is a scaling factor (set to $1$ in our experiments).
-$f_{FLCG}(\mathcal{A}\mid\mathcal{D}_{E})=\sum_{i\in\mathcal{D}}\max\left(\max_{j\in\mathcal{A}}s_{ij}-\nu\max_{k\in\mathcal{D}_{E}}s_{ik},0\right),$ (5)
-Each submodular function, when combined with our utility-based kernel, guides the selection of data subsets tailored to the specific fine-tuning stage.
-This ensures that DELIFT selects the most informative and diverse examples, maximizing the efficiency and effectiveness of fine-tuning.
-3.4 Utility Kernel as Feature Space
-Our approach utilizes the utility-based kernel as a feature space for data selection, representing a significant departure from traditional semantic similarity-based methods.
-Traditional methods often rely on sentence embeddings (SE) to capture static semantic similarities between data points.
-In contrast, our utility-based kernel measures the actual impact of examples on model performance, providing a dynamic and task-specific assessment.
-This distinction is crucial for two main reasons: 1.
-Semantic Diversity vs.
-Performance Enhancement: While SE-based methods select diverse examples solely based on semantic content, our utility-based approach selects examples that demonstrably improve model performance across various inputs; 2.
-Model-Aware Selection: The utility-based kernel is attuned to the model’s current capabilities and weaknesses, enabling the selection of data points that are most beneficial for enhancing performance on the target task.
-By integrating the utility-based kernel with the aforementioned submodular functions DELIFT tailors the data selection process to each fine-tuning stage: instruction tuning, task-specific fine-tuning, and continual learning.
-3.5 Data Subset Selection Algorithm
-To operationalize our data selection approach, we employ a **greedy algorithm** that iteratively builds the subset $\mathcal{A}$ by selecting the data point that offers the maximum marginal gain in the chosen submodular function.
+Our goal is to efficiently identify a subset of data that maximizes the performance of large language models across three fine-tuning stages: (1) Instruction Tuning, (2) Task-Specific Fine-Tuning, and (3) Continual Fine-Tuning.
+This section first introduces our novel information-theoretic pairwise utility metric (Section 3.1) and then explains how we leverage submodular optimization to achieve data-efficient selection (Section 3.2).
+Finally, we show how these components unify into a single-stop solution for all fine-tuning stages (Section 3.3).
+3.1 Pairwise Utility Metric
+Let $\mathcal{D}=\{(x_{i},y_{i})\}$ be a training set, where each $x_{i}$ is an input sequence and $y_{i}$ is the corresponding output.
+Consider two samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, and let $GT_{i}$ denote the ideal “ground truth” distribution that assigns probability 1 to the token sequence $y_{i}$ and 0 otherwise.
+We define $p(y_{i}\mid x_{i})$ as the model’s predicted probability distribution of $y_{i}$ given $x_{i}$ alone, and $p(y_{i}\mid x_{i},x_{j},y_{j})$ as the predicted distribution of $y_{i}$ when $(x_{j},y_{j})$ is also provided (e.g., as an in-context example).
+Definition (Utility Function).
+We capture the information gain of $(x_{j},y_{j})$ for predicting $(x_{i},y_{i})$ via:
+$UF_{ij}\;=\;d\bigl{(}GT_{i},\;p(y_{i}\mid x_{i})\bigr{)}\;-\;d\bigl{(}GT_{i},\;p(y_{i}\mid x_{i},x_{j},y_{j})\bigr{)},$ (1)
+where $d(\cdot,\cdot)$ is a distance between probability distributions.
+Information-Theoretic Interpretation.
+Below we state a simplified version of our main theoretical result (see Appendix A for the full proof):
+Theorem 1
+(Informal Statement) If $d(\cdot,\cdot)$ is chosen to be the Kullback-Leibler (KL) divergence, then the utility $UF_{ij}$ coincides with the (conditional) pointwise mutual information between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$.
+Formally,
+$UF_{ij}\;=\;\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}\;=\;\sum_{t=1}^{T}\log\Bigl{(}\tfrac{p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})}{p(y_{it}\mid x_{i},y_{i,<t})}\Bigr{)}.$
+Thus, $UF_{ij}$ captures how much $(x_{j},y_{j})$ informs the prediction of $y_{i}$ given $x_{i}$.
+Practical Computation.
+In practice, for numerical stability, we adopt a length-normalized Euclidean distance rather than the KL-divergence:
+$d\bigl{(}GT_{i},\;p(y_{i}\mid\cdot)\bigr{)}\,=\,\Bigl{\lVert}1-p(y_{i}\mid\cdot)\Bigr{\rVert}_{2},$
+where we only extract the probability assigned to each ground-truth token under teacher forcing.
+This preserves the spirit of the PMI-based formulation while avoiding the instability issues often encountered with near-zero probabilities in large vocabularies.
+We compute $UF_{ij}$ for all pairs $(i,j)$ once before fine-tuning.
+Although this step is $O(n^{2})$ in the dataset size, the cost is amortized because the same utility matrix can be reused for different fine-tuning stages or methods.
+3.2 Submodular Optimization for Data Selection
+After computing $UF_{ij}$, we define a kernel matrix $s_{ij}$ (e.g., set $s_{ij}=\max(UF_{ij},0)$) and utilize it in well-studied submodular functions (Fujishige, 2005; Iyer et al., 2021).
+Submodularity naturally captures diminishing returns and promotes coverage of diverse yet informative samples.
+Objectives.
+We primarily adopt three variants:
+  1. 1.
+    Facility Location (FL): $f_{FL}(\mathcal{A})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}.$
+  2. 2.
+    Facility Location Mutual Information (FLMI): $f_{FLMI}(\mathcal{A};\mathcal{D}_{T})=\sum_{i\in\mathcal{D}}\max_{j\in\mathcal{A}}s_{ij}\;+\;\eta\sum_{j\in\mathcal{A}}\max_{i\in\mathcal{D}_{T}}s_{ij}.$
+  3. 3.
+    Facility Location Conditional Gain (FLCG): $f_{FLCG}(\mathcal{A}\mid\mathcal{D}_{E})=\sum_{i\in\mathcal{D}}\max\Bigl{(}\max_{j\in\mathcal{A}}s_{ij}\;-\;\nu\,\max_{k\in\mathcal{D}_{E}}s_{ik},0\Bigr{)}.$
+FL supports coverage of $\mathcal{D}$ itself, ideal for instruction tuning; FLMI additionally ties the selection to a task-specific dataset $\mathcal{D}_{T}$; and FLCG incorporates a previously used set $\mathcal{D}_{E}$ to promote continual learning of novel examples.
 Algorithm 1 Greedy Maximization for Submodular Function
@@
 7:return $\mathcal{A}$
-This greedy algorithm ensures that each addition to the subset $\mathcal{A}$ maximizes the marginal gain in the submodular function $f$.
-By iteratively selecting the most beneficial data points according to the utility-based kernel and the specific submodular function tailored to the fine-tuning stage, DELIFT efficiently utilizes the data budget to select the most informative examples.
-The complete subset selection process involves the following steps: 1.
-Compute the Utility-Based Kernel: Calculate $UF_{ij}$ for all relevant pairs of data points in the dataset to assess their informativeness; 2.
-Select the Appropriate Submodular Function: Depending on the fine-tuning stage (instruction tuning, task-specific fine-tuning, or continual fine-tuning), choose the corresponding submodular function (FL, FLMI, or FLCG); 3.
-Apply the Greedy Maximization Algorithm: Use Algorithm 1 to iteratively build the subset $\mathcal{A}$ by selecting data points that offer the highest marginal gain according to the selected submodular function.
-By synergizing our novel utility-based kernel with submodular optimization, DELIFT achieves data-efficient fine-tuning that effectively addresses both redundancy and informativeness in the data selection process, optimizing the model’s performance across various tasks and domains.
+Subset Selection.
+To choose a subset $\mathcal{A}$ of size $k$, we apply a classic greedy heuristic (Nemhauser et al., 1978), each step picking $d^{*}=\arg\max_{d\in\mathcal{D}\setminus\mathcal{A}}[f(\mathcal{A}\cup\{d\})-f(\mathcal{A})]$ (see Algorithm 1).
+This yields a $1-\tfrac{1}{e}$ approximation factor for submodular functions, ensuring that we obtain near-optimal subsets efficiently.
+3.3 A Single-Stop Solution for All Fine-Tuning Stages
+By combining the utility-based kernel with the above submodular objectives, we obtain DELIFT, a unified framework that selects data holistically across instruction tuning, task-specific fine-tuning, and continual learning:
+  • Instruction Tuning: Use Facility Location (FL) for diverse coverage of general instructions.
+  • Task-Specific Fine-Tuning: Use FLMI to align samples with a target benchmark or domain $\mathcal{D}_{T}$.
+  • Continual Fine-Tuning: Use FLCG to incorporate new, complementary data while avoiding redundancy with existing knowledge $\mathcal{D}_{E}$.
+Despite these distinct objectives, each stage uses the same underlying pairwise utility metric to capture how one sample can enhance the model’s predictions on another.
+This consistent, model-aware metric is key to DELIFT’s strong performance across diverse tasks and model sizes.
 4 Experimental Results
-We conducted extensive experiments to evaluate the effectiveness of DELIFT across various fine-tuning scenarios, model scales, and datasets.
-This section details our experimental setup, baselines, evaluation metrics, and results analysis.
+We conduct extensive experiments to evaluate DELIFT in three fine-tuning scenarios: (1) Instruction Tuning, (2) Task-Specific Fine-Tuning, (3) Continual Fine-Tuning.
+This section outlines our experimental setup and baselines, followed by results across multiple models, datasets, and fine-tuning paradigms, along with interpretative commentary to utilize the extra space for deeper insights.
+4.1 Experimental Setup
+Models.
+We evaluate DELIFT on both base LLMs (Llama-3.2-3B, Mistral-7B-v0.1, opt-125m) and instruction-tuned LLMs (Qwen2-72B-Instruct, Phi-3-mini-128k-instruct).
+This variety covers different parameter scales (millions vs.
+billions) along with diverse adaptation strategies (ICL, QLoRA, and, for smaller models, full fine-tuning).
+This diversity in model size and training setup provides a robust test-bed for assessing DELIFT across both resource-rich and more constrained environments.
+Datasets.
+We group the datasets by the primary goal of fine-tuning, ensuring a clear mapping from the data to the corresponding submodular objective.
+In particular: 1.
+Instruction Tuning: Mix-Instruct (Jiang et al., 2023), P3 (Sanh et al., 2021).
+Both aim to enhance general instruction-following behavior, featuring a variety of task prompts and user requests.
+2.
+Task-Specific Fine-Tuning: HotpotQA (Yang et al., 2018) aligned with MMLU (Hendrycks et al., 2021), Mix-Instruct aligned with MT-Bench (Zheng et al., 2023), and Mix-Instruct aligned with GSM-8k (Cobbe et al., 2021).
+These pairings allow us to extract only the most relevant samples from a large corpus to improve performance on a specific target benchmark.
+3.
+Continual Fine-Tuning: (a) SQuAD (Rajpurkar et al., 2016) paired with HotpotQA to inject more complex, multi-hop reasoning data after simpler QA, and (b) a proprietary IBM/Government domain query rewriting dataset.
+[^1] Query rewriting transforms follow-up queries (e.g., “How much is it?”) into standalone forms (e.g., “How much is the subscription for IBM Cloud?”).
+This is essential for real-world systems where user queries often rely on prior context.
+In all cases, we fixed an approximate budget of 30% for subset selection unless otherwise noted, striking a balance between data efficiency and coverage.
+Baselines.
+In addition to a Full Data baseline, where the entire training set is used, we compare DELIFT with:
+1. Random: Uniformly selects 30% of the training set to provide a simple, model-agnostic benchmark.
+2. SelectIT (Liu et al., 2024): Generates self-reflection prompts within the LLM to rate data quality, filtering out lower-quality samples.
+3. LESS (Xia et al., 2024): Employs gradient-based influence estimates, approximated via LoRA, to identify highly impactful data points for model parameter updates.
+4. DEFT-UCS: Uses sentence embeddings to cluster the dataset for diversity; although it captures semantic variety, it lacks explicit model feedback to guide selection.
+4.
+DELIFT (SE): Operates the same submodular optimization as DELIFT but replaces our utility-based kernel with static sentence embedding similarities to highlight the benefit of DELIFT’s dynamic, model-aware approach.
+These baselines range from naive (Random) to sophisticated (gradient-based, reflection-based, or embedding-based), furnishing a thorough comparison against DELIFT.
+Metrics.
+We measure model performance across different aspects:
+1. ROUGE (Lin, 2004): Focuses on $n$-gram overlap for summarization tasks or generative text alignment.
+2. BGE (Xiao et al., 2023): Evaluates semantic similarity through the dot product of normalized sentence embeddings.
+3. LAJ (LLM-as-Judge) (Kim et al., 2023): Assigns a 1–5 rating reflecting correctness, clarity, and instruction adherence; we detail the scoring rubric and example prompts in Appendix C.
+4. Classification Accuracy: Used primarily for multiple-choice tasks like MMLU.
+By combining these metrics, we obtain both text-overlap measures (ROUGE), semantic evaluations (BGE), a holistic LLM-based score (LAJ), and a classification perspective (accuracy), offering a comprehensive view of model improvements under each fine-tuning scenario.
+4.2 Use Case 1: Instruction Tuning
+Setting.
+Instruction tuning aims to broaden a model’s capacity to follow instructions spanning diverse domains and styles.
+We adopt the Facility Location (FL) objective to maximize coverage of varied instruction types.
+4.2.1 Results on Instruction-Tuned Models.
+Tables 1 and 2 compare DELIFT with baselines on Mix-Instruct and P3, using two instruction-tuned models (Qwen2-72B-Instruct and Phi-3-mini-128k-instruct).
+Across both ICL and QLoRA, DELIFT surpasses other subset selection strategies and manages to prune up to 70% of data with minimal loss.
+Interestingly, DELIFT approaches—or occasionally exceeds—the Full Data baseline, indicating strong redundancy in instruction data.
 Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 37.87 78.92 2.98 36.36 82.55 3.02 25.76 43.34 1.42 35.50 80.46 2.58 Random 39.00 80.66 3.12 44.45 85.46 3.12 33.05 72.73 2.92 44.70 83.75 2.95 SelectIT 43.08 84.50 3.18 45.14 85.88 3.21 36.11 76.31 3.18 49.68 85.84 3.20 LESS 42.08 83.24 3.26 45.16 84.95 3.28 47.10 85.94 3.23 48.68 85.86 3.24 DELIFT (SE) 47.43 84.40 3.28 48.22 86.50 3.28 46.62 85.28 3.24 45.64 83.70 3.27 DELIFT 48.46 85.77 3.35 52.79 88.04 3.37 49.83 85.27 3.32 50.31 84.40 3.33 Full Data 58.65 88.72 3.45 65.51 92.24 3.51 55.92 88.26 3.45 74.98 93.33 3.84
-Table 1: Results on Use Case 1: MixInstruct.
+Table 1: Use Case 1: MixInstruct. Bold indicates best performance. DELIFT prunes 70% data yet stays close to Full Data.
+Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 18.03 59.13 1.54 20.15 58.38 1.78 20.10 48.66 1.36 20.64 49.17 1.39 Random 20.05 59.39 1.79 20.29 59.39 1.83 20.83 49.92 2.24 24.51 53.41 2.36 SelectIT 31.38 71.08 2.86 32.96 74.76 2.90 35.37 66.67 2.52 38.98 69.84 2.54 LESS 34.59 83.23 3.07 35.03 83.37 3.50 39.69 72.12 3.17 40.32 70.89 3.24 DELIFT (SE) 34.69 83.31 3.43 35.46 83.43 3.53 37.07 71.49 3.52 38.13 79.68 3.74 DELIFT 35.48 83.69 3.58 35.60 83.64 3.54 40.66 84.00 3.68 41.91 84.53 3.76 Full Data 36.43 84.25 3.53 35.88 76.87 3.63 42.07 85.26 3.78 44.73 87.03 3.82
+Table 2: Use Case 1: P3. DELIFT again discards most data while retaining strong performance.
+4.2.2 Evaluation on a Base Model.
+Table 3 verifies these findings on a non-instruction-tuned Llama-3.2-3B (base).
+Even without prior instruction alignment, DELIFT outperforms gradient- and embedding-based baselines by a solid margin, corroborating that the utility metric is robust to model initialization.
+Method ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ Initial 25.88 61.97 1.90 28.51 73.14 2.48 Random 28.64 77.14 2.59 38.79 79.89 2.53 SelectIT 42.67 81.54 2.67 45.87 84.67 2.61 DEFT-UCS 41.55 80.12 2.63 41.03 81.86 2.59 LESS 44.99 82.48 2.69 50.54 84.14 2.78 DELIFT (SE) 51.19 83.54 2.72 55.32 85.92 2.79 DELIFT 54.58 88.29 2.83 58.57 90.98 2.98 Full Data 55.46 92.71 3.31 61.23 95.52 3.10
+Table 3: Use Case 1: Llama-3.2-3B (base) on Mix-Instruct (30% subset).
+Takeaway (Use Case 1): The FL objective and utility-based kernel consistently prune large volumes of instruction data without undermining final performance—a strong indicator that many instruction examples are either repetitive or uninformative.
+4.3 Use Case 2: Task-Specific Fine-Tuning
+Setting.
+We refine models for specialized domains (reasoning, QA). Using Facility Location Mutual Information (FLMI), we select examples aligned with a target dataset $\mathcal{D}_{T}$.
+4.3.1 HotpotQA $\to$ MMLU & MixInstruct $\to$ MT-Bench.
+Method Qwen2 (QLoRA) Phi-3 (QLoRA) Initial 82.10 69.10 Random 79.31 65.16 SelectIT 79.13 65.24 LESS 80.35 66.72 DELIFT (SE) 80.10 66.36 DELIFT 81.70 68.70 Full Data 78.36 64.50
+Table 4: Use Case 2: HotpotQA and MMLU (5-shot) for Qwen2 and Phi-3 (classification accuracy). DELIFT outperforms Full Data by 3.34% for Qwen2 and by 4.20% for Phi-3.
+Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 44.32 74.86 2.48 47.65 77.92 2.72 39.57 69.43 2.31 42.89 72.76 2.53 Random 49.78 79.54 2.83 52.91 82.67 3.05 44.63 74.28 2.62 47.85 77.39 2.84 SelectIT 54.92 83.71 3.12 57.86 86.59 3.31 49.75 78.64 2.91 52.68 81.52 3.13 LESS 59.63 85.89 3.29 62.74 88.72 3.48 54.82 81.95 3.08 57.73 84.67 3.29 DELIFT (SE) 62.85 86.94 3.38 65.83 89.76 3.57 57.69 82.87 3.17 60.54 85.59 3.38 DELIFT 64.73 87.82 3.47 67.91 90.64 3.66 59.58 83.76 3.26 62.47 86.48 3.47 Full Data 65.89 88.65 3.55 69.72 91.53 3.74 60.76 84.59 3.34 64.31 87.42 3.55
+Table 5: Use Case 2: Mix-Instruct and MT-Bench.
 Bold indicates the best performance.
-There is a 10.44% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.27% performance percentage drop from DELIFT to the next best baseline.
-Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 18.03 59.13 1.54 20.15 58.38 1.78 20.10 48.66 1.36 20.64 49.17 1.39 Random 20.05 59.39 1.79 20.29 59.39 1.83 20.83 49.92 2.24 24.51 53.41 2.36 SelectIT 31.38 71.08 2.86 32.96 74.76 2.90 35.37 66.67 2.52 38.98 69.84 2.54 LESS 34.59 83.23 3.07 35.03 83.37 3.50 39.69 72.12 3.17 40.32 70.89 3.24 DELIFT (SE) 34.69 83.31 3.43 35.46 83.43 3.53 37.07 71.49 3.52 38.13 79.68 3.74 DELIFT 35.48 83.69 3.58 35.60 83.64 3.54 40.66 84.00 3.68 41.91 84.53 3.76 Full Data 36.43 84.25 3.53 35.88 76.87 3.63 42.07 85.26 3.78 44.73 87.03 3.82
-Table 2: Results on Use Case 1: P3.
-Bold indicates the best performance.
-There is only a 0.76% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.23% performance percentage drop from DELIFT to the next best baseline.
-4.1 Datasets and Use Cases
-We evaluated DELIFT across the three previously described fine-tuning scenarios:
-Use Case 1: Instruction Tuning We evaluated the effectiveness of DELIFT for use case 1 on two datasets: MixInstruct (Jiang et al., 2023) and P3 (Public Pool of Prompts) (Sanh et al., 2021).
-We randomly selected 21,000 train, 6,000 valid, and 3,000 test samples.
-Using the Facility Location (FL) submodular function, we aimed to select a subset of training data that was both representative and informative.
-Use Case 2: Task-Specific Fine-Tuning We evaluated DELIFT for task-specific fine-tuning using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct with MT-Bench (Zheng et al., 2023).
-We used the Facility Location Mutual Information (FLMI) submodular function to select the most informative samples from the training datasets (HotpotQA and MixInstruct) that shared relevant information with the target datasets (MMLU and MT-Bench, respectively).
-Use Case 3: Continual Fine-Tuning We evaluated DELIFT in a continual fine-tuning setting using two dataset pairs: (1) SQuAD (Rajpurkar et al., 2016) paired with HotpotQA for general question-answering, and (2) proprietary query rewriting datasets covering IBM and government domains.
-[^1] In this task, non-standalone questions –questions that require previous context to answer– must be rewritten to be standalone.
-For example, ”How much is it?” should be rewritten to ”How much is the subscription for IBM Cloud?” Such queries are common in user-agent conversations where a user asks a follow-up to an agent.
-Our goal was to integrate new knowledge efficiently while minimizing redundancy.
-We employed the Facility Location Conditional Gain (FLCG) submodular function, selecting complementary samples from the new dataset (HotpotQA and Government query rewrite) that provided additional, non-overlapping information to the existing dataset (SQuAD and IBM query rewrite).
-4.2 Experimental Setup
-Models: We evaluated DELIFT on two state-of-the-art open-source models: Phi-3-mini-128k-instruct (Abdin et al., 2024): 3.8B parameters, Qwen2-72B-Instruct (Yang et al., 2024): 72B parameters.
-These models were chosen to demonstrate effectiveness across different model scales.
-Model Qwen2 Phi-3 Method QLoRA QLoRA Initial 82.10 69.10 Random 79.31 65.16 SelectIT 79.13 65.24 LESS 80.35 66.72 DELIFT (SE) 80.10 66.36 DELIFT 81.70 68.70 Full Data 78.36 64.50
-Table 3: Results on Use Case 2: HotpotQA and MMLU (5-shot) for Qwen2 and Phi-3 models (classification accuracy).
-Bold indicates the best performance.
-For Qwen2, DELIFT outperforms Full Data by 3.34%, while for Phi-3, it improves by 4.20%.
-Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 44.32 74.86 2.48 47.65 77.92 2.72 39.57 69.43 2.31 42.89 72.76 2.53 Random 49.78 79.54 2.83 52.91 82.67 3.05 44.63 74.28 2.62 47.85 77.39 2.84 SelectIT 54.92 83.71 3.12 57.86 86.59 3.31 49.75 78.64 2.91 52.68 81.52 3.13 LESS 59.63 85.89 3.29 62.74 88.72 3.48 54.82 81.95 3.08 57.73 84.67 3.29 DELIFT (SE) 62.85 86.94 3.38 65.83 89.76 3.57 57.69 82.87 3.17 60.54 85.59 3.38 DELIFT 64.73 87.82 3.47 67.91 90.64 3.66 59.58 83.76 3.26 62.47 86.48 3.47 Full Data 65.89 88.65 3.55 69.72 91.53 3.74 60.76 84.59 3.34 64.31 87.42 3.55
-Table 4: Results on Use Case 2: MixInstruct and MT-Bench.
-Bold indicates the best performance.
-There is a 2.91% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 1.14% performance percentage drop from DELIFT to the next best baseline.
-Metrics: We use a variety of metrics to characterize performance.
-For $n$-gram word overlap we use ROUGE (Lin, 2004).
-For semantic similarity we calculate the dot product between the embeddings from the bge-large-en-v1.5 model (Xiao et al., 2023); the embeddings are normalized to unit vectors, hence the closer the dot product is to 1, the more semantically similar the vectors (the metric is referred to as ‘BGE’).
-Additionally, we use Prometheus (Kim et al., 2023), specifically the prometheus-7b-v2.0 model, as an LLM-as-a-Judge (referred to as ‘LAJ’).
-With our custom rubric outlined in Appendix B, Prometheus assigns scores in a range of 1 to 5 (higher scores indicate better performance.) Finally, we use classification accuracy to evaluate MMLU.
-Baselines: We evaluated DELIFT by comparing it against several baselines to understand its effectiveness in data selection.
-These baselines included: (1) SelectIT (Liu et al., 2024), which selects data using model feedback at the token, sentence, and model levels to identify useful samples; (2) LESS (Xia et al., 2024), which leverages LoRA approximated gradient-based influence estimation to prioritize impactful data points; (3) Random, which selects a fixed percentage (x%) of the dataset randomly, providing a benchmark for non-strategic selection; (4) DELIFT with Sentence Embedding Features (SE), which uses DELIFT but substitutes sentence embeddings as the feature space, employing a model-independent, pairwise similarity kernel instead of the utility kernel for submodular optimization; and (5) Full Data, where the entire dataset is used for fine-tuning, serving as an upper benchmark for performance.
-For In-Context Learning (ICL), the selected subsets from each baseline were used as the pool of examples, allowing us to evaluate how effectively each method supports ICL by providing relevant and informative data.
+There is a 2.91% drop from Full Data to DELIFT after pruning 70% of the data, and a 1.14% drop from DELIFT to the next best baseline.
+Empirical Findings.
+Tables 5 and 4 illustrate DELIFT’s consistent success in bridging the gap between diverse source datasets (Mix-Instruct or HotpotQA) and target benchmarks (MT-Bench, MMLU).
+In some cases (HotpotQA $\to$ MMLU), DELIFT surpasses the Full Data baseline by pruning unhelpful training points.
+4.3.2 Further Experiments on Complex Reasoning Tasks.
+To rigorously test DELIFT on more challenging datasets, we apply it to Mistral-7B-v0.1 (base) using Mix-Instruct aligned with GSM-8k, a corpus of math word problems that demand multi-step logical reasoning.
+Table 6 shows that DELIFT yields strong improvements under both ICL and QLoRA, surpassing all competing baselines.
+Even the simpler FL-only variant outperforms LESS and DEFT-UCS, highlighting the advantage of carefully selected examples when dealing with more complex reasoning tasks.
+Moreover, using FLMI for task alignment confers an additional boost, reinforcing the importance of matching the submodular objective to the dataset’s complexity and problem-solving requirements.
+Method ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ Initial 17.06 55.13 1.35 27.95 65.92 1.36 Random 19.18 56.03 1.38 33.85 81.74 1.85 SelectIT 30.27 77.23 1.45 42.29 88.17 2.49 DEFT-UCS 30.06 76.99 1.59 41.45 87.84 2.08 LESS 31.69 77.87 2.26 43.86 88.22 2.60 DELIFT (SE) 31.84 78.62 2.44 43.04 90.53 2.54 DELIFT (FL instead of FLMI) 32.30 78.54 2.54 44.62 91.04 2.63 DELIFT 33.25 79.10 2.56 46.12 92.97 2.71 Full Data 35.33 81.59 2.79 49.10 94.57 2.85
+Table 6: Use Case 2: Mistral-7B-v0.1 (base): Mix-Instruct $\to$ GSM-8k with FLMI.
+Takeaway (Use Case 2): By explicitly matching data to a target domain or benchmark, DELIFT discards irrelevant samples and—in some cases—yields better results than training on the entire dataset.
+4.4 Use Case 3: Continual Fine-Tuning
+Setting.
+A model must assimilate new data without forgetting old knowledge. We use Facility Location Conditional Gain (FLCG) to prioritize novel samples while avoiding overlaps.
+4.4.1 Examples: IBM $\to$ Government, SQuAD $\to$ HotpotQA.
+Tables 7 and 8 confirm that DELIFT mitigates redundancy. Notably, on IBM $\to$ Government, DELIFT sees less than 1% drop vs. Full Data, versus a 3–4% gap for baselines like LESS and DEFT-UCS.
 Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 44.11 70.49 2.43 48.49 80.85 2.62 40.66 58.68 1.52 43.96 69.56 2.29 Random 55.57 85.26 2.91 55.52 85.53 2.94 45.76 76.19 2.45 58.94 82.41 2.89 SelectIT 63.07 86.38 3.18 65.42 87.50 3.20 63.49 85.27 2.96 64.09 85.07 3.16 LESS 64.28 85.41 3.29 69.85 89.33 3.45 66.01 87.20 3.19 67.53 88.17 3.22 DELIFT (SE) 61.07 85.16 3.45 74.05 92.47 3.58 68.84 88.46 3.32 69.30 88.62 3.35 DELIFT 69.49 87.94 3.60 74.19 92.23 3.65 74.11 89.41 3.57 74.38 91.55 3.57 Full Data 66.08 87.84 3.65 76.83 92.63 3.74 71.23 91.10 3.52 77.12 91.10 3.64
-Table 5: Results on Use Case 3: IBM and Government.
-Bold indicates the best performance.
-There is only a 0.31% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.89% performance percentage drop from DELIFT to the next best baseline.
+Table 7: Use Case 3: IBM and Government. DELIFT effectively selects novel examples, outperforming baselines.
 Model Qwen2 Phi-3 Method ICL QLoRA ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 51.51 66.97 1.77 54.18 78.27 2.50 40.42 58.23 1.26 40.94 58.12 1.29 Random 54.38 79.12 2.57 59.23 82.02 2.66 44.29 59.45 1.33 50.29 61.52 1.60 SelectIT 58.03 83.75 2.82 63.26 84.01 2.87 47.35 74.15 2.54 56.88 80.47 2.70 LESS 67.16 85.76 2.94 69.72 86.63 3.26 60.97 81.41 2.84 61.56 81.53 2.88 DELIFT (SE) 73.75 88.01 3.26 74.84 88.79 3.30 64.44 83.95 3.03 66.35 84.77 3.14 DELIFT 76.94 90.41 3.33 77.56 89.99 3.34 66.55 84.65 3.25 67.09 85.17 3.32 Full Data 77.78 90.31 3.35 78.72 90.77 3.48 68.47 85.93 3.33 70.48 86.06 3.44
-Table 6: Results on Use Case 3: SQuAD and HotpotQA.
-Bold indicates the best performance.
-There is only a 1.94% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.78% performance percentage drop from DELIFT to the next best baseline.
-4.3 Results and Analysis
-To ensure a fair and comprehensive evaluation of DELIFT, we conducted experiments across three distinct fine-tuning scenarios: instruction tuning, task-specific fine-tuning, and continual fine-tuning.
-For all subset selection methods—including DELIFT, Random, SelectIT, LESS, and DELIFT with Sentence Embdedding Features (SE)—we consistently selected 30% of the dataset as a subset, enabling direct comparisons between methods and with the full dataset baseline (see Section 4.4 for an ablation study examining the impact of subset size).
-Use Case 1: Instruction Tuning Our first set of experiments focused on instruction tuning, a crucial task to enhancing a model’s ability to follow diverse instructions.
-As shown in Tables 1 and 2, DELIFT achieved a minimal performance drop of only 5.60% compared to using the full dataset while reducing the dataset by 70%.
-This demonstrates DELIFT’s capability to retain the most informative samples essential for instruction tuning.
-Furthermore, DELIFT outperformed other subset selection methods, achieving a 2.74% improvement and a substantial 26.21% advantage over the next best and worst-performing baselines, respectively.
-These results underscore DELIFT’s superior ability to maintain high performance with significantly reduced data, highlighting its efficacy in instruction tuning.
-Use Case 2: Task-Specific Fine-Tuning In the task-specific fine-tuning scenario, we evaluated DELIFT using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct paired with MT-Bench (Zheng et al., 2023).
-Results, presented in Tables 4 and 3, demonstrate DELIFT’s consistent and competitive performance across different task pairs.
-A particularly noteworthy outcome emerged from the HotpotQA-MMLU pair, where DELIFT not only matched but exceeded the performance of the full dataset, achieving a 5.51% improvement.
-This indicates that DELIFT’s selective approach can effectively filter out noise and focus on the most relevant and informative samples, yielding enhanced task-specific adaptation even with reduced data.
-Use Case 3: Continual Fine-Tuning The third use case examined DELIFT’s efficacy in continual fine-tuning, where models need to incorporate new information while retaining previously learned knowledge.
-As detailed in Tables 5 and 6, DELIFT demonstrated remarkable consistency, showing only a marginal 1.13% performance drop compared to using the full dataset.
-Moreover, DELIFT outperformed the second-best baseline by 3.33% and the worst baseline by 23.88%, highlighting its superiority in data selection.
-In specialized tasks such as query rewriting, DELIFT even surpassed the performance of the full dataset, suggesting that its selective approach effectively prunes noisy or irrelevant data points, thereby enhancing model performance.
-4.4 Ablation Study: Impact of Subset Size
-To assess how subset size influences DELIFT’s performance, we conducted an ablation study by varying the subset size from 5% to 100% of the full dataset across three use cases.
-The results, detailed in Appendix A and illustrated in Figure 2, show that LAJ scores generally increase with subset size.
-Utilizing the full dataset consistently yields the highest performance, highlighting the benefits of larger training sets.
-However, for methods such as DELIFT, SelectIT, and LESS, performance gains plateau or slow beyond a 50% subset size, indicating that additional data offers minimal benefits and may introduce redundancy.
-Importantly, DELIFT outperforms all baselines across subset sizes from 5% to 100%, demonstrating its robustness and effectiveness in selecting informative samples regardless of subset size.
-These findings suggest that carefully selected smaller datasets can achieve comparable performance to larger, unfiltered datasets, which is particularly valuable for resource-intensive large language models.
-4.5 Discussion
-The comprehensive results across all three use cases highlight DELIFT’s effectiveness and versatility.
-By consistently reducing data requirements by up to 70% while maintaining—and in some cases improving—performance, DELIFT addresses a critical challenge in large language model fine-tuning.
-The superior performance of DELIFT can be attributed to its novel pairwise utility metric and the use of tailored submodular functions for each fine-tuning stage.
-This approach enables DELIFT to select not only representative and diverse samples but also to reduce noise present in the full dataset.
-The ability to outperform full datasets in certain scenarios, particularly in niche tasks like query rewriting, underscores DELIFT’s capacity to distill the most relevant and informative data points.
-These findings have significant implications for the accessibility and efficiency of LLM fine-tuning.
-By dramatically reducing the amount of data required for effective fine-tuning, DELIFT paves the way for more widespread adoption and application of large language models across various domains, especially in resource-constrained environments.
-Furthermore, DELIFT’s consistent outperformance of existing data selection techniques across various fine-tuning scenarios and model scales demonstrates its robustness and broad applicability, making it a valuable tool for researchers and practitioners alike.
-In conclusion, our experimental results firmly establish DELIFT as a powerful and efficient method for data selection in LLM fine-tuning.
-By addressing the critical challenge of optimal data selection, DELIFT not only enhances the efficiency of model training but also opens new possibilities for fine-tuning large language models in domains where data or computational resources may be limited.
+Table 8: Use Case 3: SQuAD and HotpotQA.
+Bold indicates best performance.
+There is only a 1.94% drop from Full Data to DELIFT after pruning 70% of the data, and a 2.78% drop from DELIFT to the next best baseline.
+Takeaway (Use Case 3): By prioritizing complementary rather than redundant data, DELIFT preserves previous capabilities and focuses on new insights.
+4.5 Further Comparisons and Ablations
+4.5.1 Full Fine-Tuning vs. QLoRA.
+We also validated DELIFT under full fine-tuning on a smaller opt-125m (Table 9).
+Results show consistent gains for DELIFT, indicating that the utility-based selection approach is independent of the particular fine-tuning methodology.
+Method QLoRA Full Fine-Tuning ROUGE BGE LAJ ROUGE BGE LAJ Initial 9.04 40.50 1.19 9.57 40.86 1.14 Random 12.55 46.99 1.25 13.07 47.91 1.30 SelectIT 14.78 49.80 1.26 15.35 50.42 1.28 DEFT-UCS 15.12 50.29 1.39 15.16 50.29 1.39 LESS 15.52 50.70 1.38 16.02 51.30 1.40 DELIFT (SE) 18.81 55.02 1.35 17.06 55.93 1.38 DELIFT 19.72 57.98 1.43 19.87 58.11 1.45 Full Data 20.39 60.39 1.95 21.64 61.70 2.05
+Table 9: opt-125m on Mix-Instruct, comparing QLoRA vs. full fine-tuning.
+4.5.2 Comparing Submodular Objectives.
+Although FL alone can beat naive baselines in specialized or incremental settings (see Table 6), the specialized objectives (FLMI for domain tasks, FLCG for continual updates) yield stronger alignment.
+This underscores the importance of matching the submodular objective to the fine-tuning stage.
+4.5.3 Ablation on Subset Size.
+Beyond consistently using 30% of the data in our main experiments, we investigated how varying the subset size influences performance.
+We tested budgets ranging from as little as 5% up to 50% of the original training set (in increments of 10%).
+As shown in Figure 2, performance under all methods generally improves with larger subsets but exhibits diminishing returns beyond 30–40%.
+Notably, DELIFT consistently yields higher LAJ scores than other baselines at every subset size, demonstrating its robustness even under very aggressive pruning (e.g., 5%).
+These findings reinforce that substantial reductions in training data are possible without significant performance degradation, and that DELIFT identifies highly informative samples more effectively than competing methods.
+4.6 Key Observations
+Our experiments reveal four consistent patterns across datasets, model scales, and adaptation methods:
+  1. 1.
+    Utility-based kernel outperforms static or gradient-based methods, indicating that measuring how one sample improves predictions on another is a stronger signal than mere embeddings or approximate gradients.
+  2. 2.
+    Stage-specific objectives (FL, FLMI, FLCG) are crucial for addressing the unique needs of instruction tuning (diverse coverage), task-specific alignment, and continual complementarity.
+  3. 3.
+    Significant pruning (up to 70%) is routinely possible, often with minimal or zero performance deterioration, suggesting vast redundancy in large corpora.
+  4. 4.
+    Method-agnostic gains, as DELIFT consistently improves results under ICL, QLoRA, and even full fine-tuning on smaller models.
+In the next section, we summarize these findings, discuss limitations such as the $O(n^{2})$ cost of utility computation, and propose directions for future work, including adaptive selection during training, extensions to multimodal data, and fairness considerations.
 5 Conclusion, Limitations, and Future Work
@@
 6 Acknowledgement
-This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS240550 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
+A part of this work used the Delta system at the National Center for Supercomputing Applications through allocation CIS240550 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
 Appendix
-Appendix A Subset Size Comparison
+Appendix A Theoretical Foundations and Connections Between the Utility Metric and Information Theory
+Theorem 2
+Let $y_{i}=(y_{i1},y_{i2},\dots,y_{iT})$ be a sequence of tokens with ground truth distribution $GT_{i}$, where $GT_{i}$ assigns probability 1 to the sequence $y_{i}$ and 0 to all other sequences.
+Let $p(y_{i}\mid x_{i})$ be the predicted probability of $y_{i}$ given input $x_{i}$, and $p(y_{i}\mid x_{i},x_{j},y_{j})$ be the predicted probability of $y_{i}$ given $x_{i}$ and an in-context example $(x_{j},y_{j})$.
+Define the utility metric $UF_{ij}$ using a general distance metric $d(\cdot,\cdot)$ between probability distributions:
+$UF_{ij}=d(GT_{i},\,p(y_{i}\mid x_{i}))-d(GT_{i},\,p(y_{i}\mid x_{i},x_{j},y_{j})).$
+Claim: When the distance metric $d(\cdot,\cdot)$ is the Kullback-Leibler divergence $D_{\text{KL}}$, the utility metric $UF_{ij}$ is equal to the pointwise mutual information (PMI) between the sequence $y_{i}$ and the in-context example $(x_{j},y_{j})$ conditioned on $x_{i}$:
+$UF_{ij}=\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
+Furthermore, $UF_{ij}$ can be expressed as the sum of conditional PMI over the tokens in $y_{i}$:
+$UF_{ij}=\sum_{t=1}^{T}\operatorname{PMI}\big{(}y_{it};\,x_{j},y_{j}\mid x_{i},y_{i,<t}\big{)},$
+where $y_{i,<t}=(y_{i1},y_{i2},\dots,y_{i(t-1)})$ denotes the sequence of previous tokens up to position $t-1$.
+Proof:
+1. Computing KL-Divergence Between Ground Truth and Predicted Distributions:
+Since $GT_{i}$ assigns probability 1 to the specific sequence $y_{i}$, the KL-divergence simplifies as follows:
+$d(GT_{i},\,p(y_{i}\mid\cdot))=D_{\text{KL}}(GT_{i}\parallel p(y_{i}\mid\cdot))=-\log p(y_{i}\mid\cdot),$
+because the KL-divergence between a one-hot distribution and any other distribution reduces to the negative log-probability of the assigned event.
+2. Computing the Utility Metric $UF_{ij}$:
+The utility metric becomes:
+$UF_{ij}$ $=-\log p(y_{i}\mid x_{i})+\log p(y_{i}\mid x_{i},x_{j},y_{j})$
+$=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
+3. Expressing $UF_{ij}$ as Pointwise Mutual Information:
+The conditional pointwise mutual information between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$ is defined as:
+$\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})=\log\frac{p(y_{i},x_{j},y_{j}\mid x_{i})}{p(y_{i}\mid x_{i})\,p(x_{j},y_{j}\mid x_{i})}.$
+Using the chain rule:
+$p(y_{i},x_{j},y_{j}\mid x_{i})=p(y_{i}\mid x_{i},x_{j},y_{j})\,p(x_{j},y_{j}\mid x_{i}).$
+Substituting back:
+$\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})$ $=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})\,p(x_{j},y_{j}\mid x_{i})}{p(y_{i}\mid x_{i})\,p(x_{j},y_{j}\mid x_{i})}$
+$=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
+Therefore:
+$UF_{ij}=\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i}).$
+4. Expanding $UF_{ij}$ as Sum of Conditional PMI Terms:
+We expand $p(y_{i}\mid\cdot)$ using the chain rule:
+$p(y_{i}\mid\cdot)=\prod_{t=1}^{T}p(y_{it}\mid\cdot,y_{i,<t}),$
+where $y_{i,<t}$ is the sequence of previous tokens up to time $t-1$.
+Substituting back into $UF_{ij}$:
+$UF_{ij}$ $=\log\frac{\prod_{t=1}^{T}p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})}{\prod_{t=1}^{T}p(y_{it}\mid x_{i},y_{i,<t})}$
+$=\sum_{t=1}^{T}\left[\log p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})-\log p(y_{it}\mid x_{i},y_{i,<t})\right]$
+$=\sum_{t=1}^{T}\operatorname{PMI}\big{(}y_{it};\,x_{j},y_{j}\mid x_{i},y_{i,<t}\big{)}.$
+This shows that $UF_{ij}$ is the sum of the conditional PMI of each token $y_{it}$ with $(x_{j},y_{j})$ given $x_{i}$ and the previous tokens.
+Conclusion:
+When $d(\cdot,\cdot)=D_{\text{KL}}$, the utility metric $UF_{ij}$ precisely equals the conditional PMI between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$.
+A.1 Why Euclidean Distance is Preferred Over KL-Divergence for Subset Selection
+The effectiveness of subset selection algorithms, including facility location functions, depends critically on the properties of the chosen distance metric d(·,·).
+Euclidean distance offers several key advantages over KL-divergence for this purpose:
+  1. 1.
+    Mathematical Properties
+    • Euclidean distance is non-negative, finite, and symmetric (d(a,b) = d(b,a))
+    • KL-divergence can be infinite or undefined with zero probabilities and lacks symmetry $D_{KL}(P\parallel Q)\neq D_{KL}(Q\parallel P)$
+  2. 2.
+    Computational Advantages
+    • Euclidean distance uses simple arithmetic operations (subtraction, squares, square roots)
+    • KL-divergence requires more complex logarithmic calculations and division operations
+  3. 3.
+    Robustness in Practice
+    • Euclidean distance handles zero probabilities gracefully
+    • KL-divergence becomes undefined with zero probabilities, which occur frequently in real data
+Impact on Subset Selection: The facility location function requires positive, finite similarity measures to model coverage effects accurately.
+Euclidean distance satisfies these requirements, while KL-divergence’s potential negative or infinite values can disrupt optimization.
+Conclusion: While KL-divergence offers theoretical connections to mutual information, Euclidean distance provides:
+  • Guaranteed positive and finite utility metrics
+  • Superior computational efficiency
+  • Better numerical stability
+These practical advantages make Euclidean distance the preferred choice for computing the utility metric $UF_{ij}$ in subset selection algorithms.
+Appendix B Subset Size Comparison
 To assess how subset size influences the performance of DELIFT, we performed an ablation study by varying the subset size from 5% to 100% (specifically 5%, 15%, 30%, 50%, 100%) of the entire dataset across three distinct use cases.
 Figure 2 illustrates the performance metric LAJ as a function of subset size for each fine-tuning scenario.
-A.1 General Observations
+B.1 General Observations
   • Performance Increases with Subset Size: Across all methods, LAJ scores generally improve as the subset size increases.
@@
 This suggests that additional data beyond this point offers minimal benefits and may introduce redundancy.
-A.2 Detailed Analysis of Methods
-A.2.1 Initial vs. Random Selection
+B.2 Detailed Analysis of Methods
+B.2.1 Initial vs. Random Selection
   • Initial Baseline: Consistently records the lowest scores across all subset sizes, indicating that models without data-informed selection struggle to generate quality responses.
@@
 This lack of significant improvement highlights that uninformed data selection does not substantially enhance model quality.
-A.2.2 SelectIT and LESS Methods
+B.2.2 SelectIT and LESS Methods
   • LESS: Demonstrates a strong upward trend, particularly when subset sizes increase from 15% to 50%.
@@
 For subset sizes above 50%, SelectIT approaches the performance of DELIFT, suggesting its heuristic-driven selection becomes more effective with more data.
-A.2.3 DELIFT Variants
+B.2.3 DELIFT Variants
   • DELIFT vs.
@@
 In ICL fine-tuning scenarios, a 30% subset size with DELIFT nearly matches Full Data performance, demonstrating its efficiency in data reduction without significant loss in performance.
-A.3 Impact on Different Fine-Tuning Scenarios
+B.3 Impact on Different Fine-Tuning Scenarios
   • ICL vs.
@@
 a)Refer to caption b)Refer to caption c)Refer to caption d)Refer to caption e)Refer to caption f)Refer to caption
-Appendix B Prometheus Rubric
+Appendix C Prometheus Rubric
 The Prometheus model served as an LLM-as-a-Judge to evaluate response quality from different data selection methods.
@@
 Table 10: General Prometheus Rubric
-B.1 Usage Notes
+C.1 Usage Notes
   • Each response is evaluated independently based on the criteria above.
@@
   • Final LAJ scores are obtained by averaging the scores across all criteria.
-Appendix C LLM-as-Judges Scores
-In Tables 8 and 9, we show the distribution of Prometheus scores on one particular setting: Use Case 1, MixInstruct training and MixInstruct validation sets on the Qwen2-72B-Instruct model.
-These figures make clear that the average LGA scores computed in Tables 1-6 are true averages of a distribution of scores, not averages of a combination of just 1’s and 5’s.
+Appendix D LLM-as-Judges Scores
+In Tables 11 and 12, we show the distribution of Prometheus scores on one particular setting: Use Case 1, MixInstruct training and MixInstruct validation sets on the Qwen2-72B-Instruct model.
+These figures make clear that the average LGA scores computed in Tables 1-8 are true averages of a distribution of scores, not averages of a combination of just 1’s and 5’s.
 ICL QLoRA Initial [Uncaptioned image] [Uncaptioned image] Random [Uncaptioned image] [Uncaptioned image] SelectIT [Uncaptioned image] [Uncaptioned image]
@@
 The corresponding table is Table 1.
-C.1 Interpretation of Score Distributions
-C.1.1 Overall Trends
+D.1 Interpretation of Score Distributions
+D.1.1 Overall Trends
   • Score Variability: There is significant variability in score distributions across different methods.
@@
 This highlights their effectiveness in selecting informative and useful data for fine-tuning.
-C.1.2 Method-Specific Observations
+D.1.2 Method-Specific Observations
   • Initial and Random Methods: Both methods have lower mean scores (around 3.0 to 3.2) with wide spreads, suggesting inconsistent and generally lower-quality responses.
@@
     • DELIFT: Achieves the highest average scores (3.35 for ICL and 3.37 for QLoRA), outperforming all other methods and indicating its superior utility-based kernel and submodular optimization.
-C.1.3 Comparison with Full Data
+D.1.3 Comparison with Full Data
   • DELIFT vs.
@@
 DELIFT offers a more efficient alternative without major sacrifices in performance.
-Appendix D Limitations
+Appendix E Limitations
   • Dependence on Initial Data Quality: DELIFT’s effectiveness relies on the diversity and quality of the initial dataset.
@@
   • Bias Amplification Risks: The subset selection process may unintentionally amplify existing biases within the data, necessitating careful mitigation strategies.
-Appendix E Future Work
+Appendix F Future Work
   • Integration with Data Augmentation: Combining DELIFT with data augmentation techniques could further enhance the robustness and diversity of selected subsets.
@@
 Our ongoing efforts aim to ensure that DELIFT contributes to responsible and equitable AI development while maximizing efficiency.
-Appendix F Code and Data Availability
-To facilitate reproducibility and further research, we will make the DELIFT implementation and the datasets used in our experiments publicly available upon publication.
+Appendix G Code and Data Availability
+To facilitate reproducibility and further research, we will make the DELIFT implementation and the datasets used in our experiments publicly available.
 Interested researchers can access these resources through the following repository: https://anonymous.4open.science/r/optimizing-data-selection-0CD0.
-Appendix G Hyperparameter Settings
+Appendix H Hyperparameter Settings
 Consistent hyperparameter settings were maintained across all experiments to ensure reproducibility:
@@
   • Teacher Forcing Technique: Applied during utility metric computation to ensure reliable prediction accuracy measurement.
+$\operatorname{corr}_{AB}=\sum_{i\in B}p(y_{i}|x_{i})_{M_{A}}-p(y_{i}|x_{i})_{BM}$ (2)